### v4

Abstract
At present, in lattice-based linearly homomorphic signature schemes, especially under the standard model, there are very few schemes with tight security. This paper constructs the first lattice-based linearly homomorphic signature scheme that achieves tight security against existential unforgeability under chosen-message attacks (EUF-CMA) in the standard model. Furthermore, among existing schemes, the scheme proposed in this paper also offers certain advantages in terms of public key size, signature length, and computational cost.
keywords:
lattice, linearly homomorphic signature, tightly secure , standard model
1 Introduction
Homomorphic signatures refer to a type of signature that allows any entity to perform homomorphic operations on authenticated data, enabling the generation of new data. The entity can then obtain a valid signature for the new data without using the signing private key. Since homomorphic signatures were put forward, they have attracted more and more attention. The unique properties of homomorphic signatures provide a vast theoretical research space and significant scientific research value, playing an important role in many practical applications. For example, they can offer effective solutions in fields such as network coding[2, 3, 4] and cloud computing[5, 6].
The concept of homomorphic signature schemes was first introduced by R. Rivest [7] in a presentation at Cambridge University in 2000. Shortly thereafter, in 2002, Johnson et al. [8] provided the formal definition and security analysis of homomorphic signatures. Since then, various types of homomorphic signature schemes have emerged. When classifying homomorphic signature schemes based on the homomorphic operation functions, they can be categorized into linear homomorphic signatures, polynomial function homomorphic signatures, and fully homomorphic signatures.
In 2007, Zhao et al. proposed the first linearly homomorphic signature scheme [2], which allows arbitrary linear combinations of signed data, facilitating the verification of message integrity and effectively preventing contamination attacks on applications based on network coding. In 2009, Boneh et al. [9] provided the formal definition and security model for linearly homomorphic signature schemes and constructed a linearly homomorphic signature scheme for packet authentication in network coding protocols based on bilinear groups. Subsequently, further improvements were made in terms of efficiency, security, and privacy protection, leading to the proposal of numerous efficient and practical linearly homomorphic signature schemes. Early research primarily based linearly homomorphic signatures on number-theoretic assumptions, such as the discrete logarithm problem or the RSA problem [9, 10, 11, 12, 13]. These schemes are efficient but must be designed over $\mathbb{F}_{p}$ (where $p$ is a large prime) to ensure the difficulty of the RSA or discrete logarithm problem, and they cannot resist quantum computing attacks.
To resist quantum attacks, Boneh and Freeman[14] first proposed a linearly homomorphic signature scheme over $\mathbb{F}_{2}$ (the binary field) at the PKC 2011 conference, with its security based on the $k$-SIS problem ($k$-small integer solution problem). Subsequently, at the EUROCRYPT 2011 conference, Boneh and Freeman[15] introduced another linearly homomorphic signature scheme, whose security relies on the hardness of the SIS problem, and proposed a homomorphic signature scheme for polynomial functions, with security based on ideal lattices. In 2013, Wang et al.[16] improved the first linearly homomorphic signature scheme based on lattices, resulting in a new linearly homomorphic signature whose security is directly based on the standard SIS problem. All these schemes[14, 15, 16] have been proven secure in the random oracle model.
However, in reality, the ideal random oracle model is difficult to achieve. Chen et al. combined bonsai tree techniques[17] with the intersection method of two integer lattices proposed in [14], and for the first time constructed a lattice-based linearly homomorphic signature scheme over small fields in the standard model[18]. Building on Chen’s work, Lin et al. constructed two lattice-based linearly homomorphic signature schemes in the standard model by utilizing full-rank differential hash functions and linearly homomorphic chameleon hash functions[19]. At the ASIACRYPT 2016 conference, Boyen and Li[20] raised one of the open questions: how to construct a tightly secure short homomorphic signature scheme based on lattices. In 2024, Chen[21] constructed the first lattice-based, short, and tightly secure linear homomorphic signature scheme. This work builds on [18] by utilizing pseudo-random functions and the key-homomorphic algorithm Eval [22, 23]. However, regrettably, this scheme does not attain tight security within the framework of the standard EUF - CMA (Existential Unforgeability under Chosen - Message Attacks) security model.
Our Contribution. In the standard model, the only construction of a tight secure and short linearly homomorphic signature scheme based on lattices is found in reference [21]. Although [21] features shorter signatures, it has a complex key generation process (requiring the random selection of $\mathcal{O}(s+5)$ matrices and $\mathcal{O}(k)$ vectors for the public key, while the private key consists of one matrix and one vector). Furthermore, it employs the Bonsai Tree technique proposed by Cash et al.[17], which results in the final signature dimension being doubled.
Compared with [21] under the same parameters, the linearly homomorphic signature scheme we construct requires only $\mathcal{O}(1)$ randomly selected matrices and $\mathcal{O}(k)$ vectors for the public key. The private key, in turn, consists of only one matrix. Moreover, the size of our signature is half the size of that in [21].
In terms of security (as defined in Section 2), [21] proved that their scheme is tightly secure under U-ST-SCMA, but it fails to achieve tight security under the standard EUF-CMA definition. In contrast, our scheme can be proven to be tightly secure under EUF-CMA and also retains tight security under U-ST-SCMA. Therefore, the scheme we propose is the optimal lattice-based linearly homomorphic signature scheme with tight security in the standard model. A detailed comparison is provided in Table 1.
Table 1: Comparison of the [21]
Schemes Public key Signature Private key Security
[21] $\mathcal{O}(s+5)\left\|\mathbb{Z}_{q}^{h\times n}\right\|+\mathcal{O}(k)\left\|\mathbb{Z}_{q}^{h}\right\|$ $\left\|\mathbb{Z}^{2n}\right\|$ $\left\|\mathbb{Z}^{n\times n}\right\|+\left\|\mathbb{Z}_{2}^{s}\right\|$ U-ST-SCMA
Ours $\mathcal{O}(1)\left\|\mathbb{Z}_{q}^{h\times n}\right\|+\mathcal{O}(k)\left\|\mathbb{Z}_{q}^{h}\right\|$ $\left\|\mathbb{Z}^{n}\right\|$ $\left\|\mathbb{Z}^{n\times n}\right\|$ EUF-CMA/U-ST-SCMA
Our Techniques. First, in the generation of the public key. [18, 21] mainly draw on the lattice basis delegation technique algorithm ExtBasis [17] proposed by Cash et al. This technique requires additional public matrices to generate new lattices and include these matrices as part of the public key. This not only results in the dimensionality of the newly generated lattices being larger than that of the original lattices but also leads to an increase in the dimensionality of the signatures generated through trapdoor sampling from the new lattices.
In 2010, Agrawal et al. noticed the defect in this aspect of the lattice basis delegation technique proposed by Cash et al[17]. For this reason, Agrawal et al. designed a new lattice basis delegation technique algorithm $\textbf{BasisDel}(A,R,T_{A},s)$[24] in which the dimension can remain unchanged during the lattice basis delegation process. Here, $A\in\mathbb{Z}_{q}^{h\times n}$ of rank $h$, $R\in\mathbb{Z}_{q}^{n\times n}$ is an invertible matrix, $T_{A}$ is the basis of $\Lambda_{q}^{\bot}(A)$, and $s$ is a positive real number.
Inspired by the algorithm BasisDel proposed by Agrawal et al., we construct a new algorithm, NewBasis$(A,H,T_{A})$, for generating trapdoors of new lattices . This algorithm is a special case of the BasisDel$(A,R,T_{A},s)$ algorithm, where the matrix $R$ is replaced by a diagonal matrix $H$ with diagonal entries equal to ±$\pm 1$, which is evidently an orthogonal matrix. Therefore, by means of the algorithm NewBasis, the public key generated by the scheme we constructed will not involve a large number of public matrices. Moreover, the signatures obtained through preimage sampling using the trapdoor generated by this algorithm have also significantly reduced in size. As a result, the signature efficiency has been greatly improved.
Secondly, regarding the generation of signatures. First, in the processing of the tag $\tau$, we map the $n$-dimensional bit string $\tau$ one-to-one to a diagonal matrix $H_{\tau}$ with diagonal elements all equal to ±$\pm 1$. Therefore, through the trapdoor algorithm TrapGen$(q,h,n)$ [25] and the algorithm NewBasis$(A,H,T_{A})$, a new trapdoor $T_{B_{\tau}}$ will be obtained. $T_{B_{\tau}}$ is the basis of the lattice $\Lambda_{q}^{\bot}(B_{\tau})$, where $B_{\tau}=AH_{\tau}^{\top}$. Finally, we will encode the message through some random vectors (which are part of the public key) to generate a vector. Subsequently, we will use the preimage sampling algorithm on the encoded vector to generate the signature of the message.
Paper Organization. In Section 2, we introduce the definition of linear homomorphic signatures and their security model. In Section 3, we present some commonly used notations and preliminary knowledge. In Section 4, we propose our linear homomorphic signature scheme and prove its correctness. In Section 5, we demonstrate that the linear homomorphic signature scheme proposed in this paper satisfies both security and privacy. The final section provides a brief summary of the paper and discusses some open problems.
2 Linearly homomorphic signature: definition and security model
Before introducing the linearly homomorphic signature, we define a multi-data homomorphic signature [21, 26, 27]. Different from the classical definition of digital signatures, this homomorphic signature consists of five polynomial-time algorithms (Setup, KeyGen, Sign, Verify, Evaluate), allowing the signer to sign multiple different data sets (each data set corresponds to a tag), and the size of each data set can be different. It is also assumed that the verifier knows the tag of the data set over which they want to verify the computation. The definition of the multi-data homomorphic signature is as follows.
Definition 1.
(Multi-data homomorphic signature, [21, 26, 27]) A multi-data homomorphic signature scheme is a tuple of probabilistic, polynomial-time algorithms (Setup, KeyGen, Sign, Verify, Evaluate) as follows:
-Setup$(1^{n},k)$. The algorithm takes as input a security parameter $n$ and and a maximum data set size $k$, and outputs a public parameter $pp$. The public parameter $pp$ defines a message space $\mathcal{M}$, a signature space $\Sigma$, and a set $\mathcal{F}$ of admissible functions $f:\mathcal{M}^{k}\longrightarrow\mathcal{M}.$
-KeyGen$(pp)$. Taking the public parameter $pp$ as input, this algorithm outputs a pair of public key and private key $(pk,sk)$
-Sign$(sk,\tau,\mathbf{m},i)$. Takes a secret key $sk$, a tag $\tau\in\{0,1\}^{n}$, a message $\mathbf{m}\in\mathcal{M}$ and index $i\in\{1,...,k\}$, and output a signature $\sigma\in\Sigma$.
-Verify$(pk,\tau,\mathbf{m},\sigma,f)$. Takes a public key $pk$, a tag $\tau\in\{0,1\}^{n}$, a message $\mathbf{m}\in\mathcal{M}$ a signature $\sigma\in\Sigma$, and a function $f\in\mathcal{F}$, and outputs either 0 (reject) or 1 (accept).
-Evaluate$(pk,\tau,f,\vec{\sigma})$. Takes a public key $pk$, a tag $\tau\in\{0,1\}^{n}$, a function $f\in\mathcal{F}$, and a tuple of signatures $\vec{\sigma}\in\Sigma^{k}$, and outputs a signature $\sigma^{\prime}\in\Sigma$.
Let $\pi_{i}:\mathcal{M}^{k}\longrightarrow\mathcal{M}$ be the function $\pi_{i}(\mathbf{m}_{1},...,\mathbf{m}_{k})=\mathbf{m}_{i}$ that projects onto the $i$-th component. We require that $\pi_{1},...,\pi_{k}\in\mathcal{F}$ for all $pk$ output by Setup$(1^{n},k)$.
For correctness, we require that for each $(pk,sk)$ output by Setup$(1^{n},k)$, it holds that:
  1. (1)
    For all tags $\tau\in\{0,1\}^{n}$, all $\mathbf{m}\in\mathcal{M}$, and all $i\in\{1,...,k\}$, if $\sigma\longleftarrow\textbf{Sign}(sk,\tau,\mathbf{m},i)$, then
    $\textbf{Verify}(pk,\tau,\mathbf{m},\sigma,\pi_{i})=1.$
  2. (2)
    For all tags $\tau\in\{0,1\}^{n}$, all $\overrightarrow{\mathbf{m}}=(\mathbf{m}_{1},...,\mathbf{m}_{k})\in\mathcal{M}^{k}$, and all $f\in\mathcal{F}$, if $\sigma_{i}\longleftarrow\textbf{Sign}(sk,\tau,\mathbf{m},i)$, for   $i\in\{1,...,k\}$, then
    $\textbf{Verify}(pk,\tau,f(\overrightarrow{\mathbf{m}}),\textbf{Evaluate}(pk,\tau,f,(\sigma_{1},...,\sigma_{k})))=1.$
We say that a signature scheme as above is multi-data $\mathcal{F}$-homomorphic.
The definition of linearly homomorphic signatures is given below.
Definition 2.
([21]) In the above definition of a multi-data homomorphic signature scheme, if the message space $\mathcal{M}$ consists of $n$-dimensional vectors over a ring $\mathbf{R}$, and the set of admissible functions $\mathcal{F}$ consists of $\mathbf{R}$-linear function from $(\mathbf{R^{n}})^{k}$ to $\mathbf{R}$, we refer to it as the multi-data linearly homomorphic signature scheme. A linear function $f(\mathbf{v}_{1},...,\mathbf{v}_{k})=\sum_{i=1}^{i=k}c_{i}\mathbf{v}_{i}$ is represented as the vector $(c_{1},...,c_{k})\in\mathbf{R}^{k}$.
Referring to [15, 19], the security model of linearly homomorphic signature is given below. (Note: The signature scheme definition provided in [15, 19] is essentially no different from Definition 2)
Definition 3.
(EUF-CMA, [15, 19]) A linearly homomorphic signature $\mathcal{LS}$=(Setup, Sign, Verify, Evaluate) satisfies existential unforgeability under chosen-message attacks (EUF-CMA) if for all $k$ the advantage of any probabilistic, polynomial-time adversary $\mathcal{A}$ in the following game is negligible in the security parameter $n$:
[…]
Queries: Proceeding adaptively, $\mathcal{A}$ specifies a sequence of data sets $\overrightarrow{\mathbf{m}}_{i}\in M^{k}$. For each $i$, the challenger choose $\tau_{i}$ uniformly from $\{0,1\}^{n}$ and gives to $\mathcal{A}$ the tag $\tau_{i}$ and the signature $\sigma_{i,j}=\textbf{Sign}(sk,\tau_{i},\mathbf{m}_{ij},j)$ for $j=1,...,k$.
Output: $\mathcal{A}$ output a tag $\tau^{*}\in\{0,1\}^{n}$, a message $\mathbf{m}^{*}\in\mathcal{M}$, a linear function $f\in\mathcal{F}$, and a signature $\sigma^{*}\in\Sigma$.
The adversary wins if $\textbf{Verify}(pk,\tau^{*},\mathbf{m}^{*},\sigma^{*},f)=1$ and either
  1. (1)
    $\tau^{*}\neq\tau_{i}$ for all $i$ (I-type forgery), or
  2. (2)
    $\tau^{*}=\tau_{i}$ for some $i$ but $\mathbf{m}^{*}\neq f(\overrightarrow{\mathbf{m}}_{i})$ (II-type forgery).
The advantage of $\mathcal{A}$ is the probability that $\mathcal{A}$ wins the security game.
In [19], an even weaker security model, EUF-naCMA, is also defined: it requires that adversary $\mathcal{A}$ submit a sequence of data sets that it wants to be signed to the challenger before seeing the public key.
In [21], the purpose of Chen is to construct a tightly secure linearly homomorphic signature Scheme. However, in practice, it cannot satisfy tight security under EUF-CMA. Therefore, inspired by [28, 29], they proposed unforgeability under the selective-tag static chosen-message-attack (U-ST-SCMA), and achieved tight security under this security model. At present, it is unclear whether there is any connection between the concept of U-ST-SCMA and the standard concept of EUF-CMA. This is one of the open questions for future research. The linearly homomorphic signature proposed in this paper also satisfies tight security under U-ST-SCMA. Below is the definition of the U-ST-SCMA security model.
Definition 4.
(U-ST-SCMA, [21]) For a linearly homomorphic signature scheme $\mathcal{LS}$=(Setup, Sign, Verify, Evaluate), we consider the following game:
Initialization: (select tag and data phase): Let $n_{s}$ be the number of the data sets . The challenger uniformly samples $(\tau_{i}\stackrel{{\scriptstyle\$}}{{\leftarrow}}\{0,1\}^{n})_{i\in\{1,...,n_{s}\}}$ and sends them to the adversary $\mathcal{A}$. Then, $\mathcal{A}$ specifies a sequence of data sets $(\overrightarrow{\mathbf{m}}_{i}=(\mathbf{m}_{i1},...,\mathbf{m}_{ik})\in\mathcal{M}^{k})_{i\in\{1,...,n_{s}\}}$ and a target tag $\tau_{g}\in\{\tau_{1},...,\tau_{n_{s}}\}$, and returns them.
[…]
Queries: The challenger computes $(\sigma_{ij}{\leftarrow}\textbf{Sign}(sk,\tau_{i},\mathbf{m}_{ij},j))_{i\in\{1,...,n_{s}\},j\in\{1,...,k\}}$, and gives the signatures $(\sigma_{ij})_{i\in\{1,...,n_{s}\},j\in\{1,...,k\}}$ to $\mathcal{A}$.
Output: $\mathcal{A}$ outputs a tag $\tau^{*}\in\{0,1\}^{n}$, a message $\mathbf{m}^{*}\in\mathcal{M}$, a linear function $f\in\mathcal{F}$, and a signature $\sigma^{*}\in\Sigma$. Then The adversary wins if $\textbf{Verify}(pk,\tau^{*},\mathbf{m}^{*},\sigma^{*},f)=1$ and either
  1. (1)
    $\tau^{*}\neq\tau_{i}$ for all $i$ (I-type forgery), or
  2. (2)
    $\tau^{*}=\tau_{g}$ but $\mathbf{m}^{*}\neq f(\overrightarrow{\mathbf{m}}_{g})$ (II-type forgery).
In homomorphic signatures, privacy also needs to be considered. Below, we provide the definition of weakly context hiding[15].
Definition 5.
(Weakly context hiding, [15]) A linearly homomorphic signature scheme $\mathcal{LS}$=(Setup, Sign, Verify, Evaluate) is weakly context hiding if for all $k$, the advantage of any probabilistic, polynomial-time adversary $\mathcal{A}$ in the following game is negligible in the security parameter $n$:
[…]
Challenge: $\mathcal{A}$ outputs $(\overrightarrow{\mathbf{m}}_{0},\overrightarrow{\mathbf{m}}_{1},f_{1},...,f_{s})$ with $\overrightarrow{\mathbf{m}}_{0},\overrightarrow{\mathbf{m}}_{1}\in\mathcal{M}^{k}$. The function $f_{1},...,f_{s}\in\mathcal{F}$ and satisfy
$f_{i}(\overrightarrow{\mathbf{m}}_{0})=f_{i}(\overrightarrow{\mathbf{m}}_{1})\quad\text{for all}\quad i=1,...,s.$
In response, the challenger generates a random bit $b\in\{0,1\}$ and a random tag $\tau\in\{0,1\}^{n}$. It signs the message in $\overrightarrow{\mathbf{m}}_{b}$ using the tag $\tau$ to obtain a vector $\overrightarrow{\sigma}$ of $k$ signatures. Next, for $i=1,...,s$ the challenger computes a signature $\sigma_{i}=\textbf{Evaluate}(pk,\tau,f_{i},\overrightarrow{\sigma})$ of $f_{i}(\overrightarrow{\mathbf{m}}_{b})$. It sends the $(\tau,\sigma_{1},...,\sigma_{s})$ to $\mathcal{A}$. Note that the functions $f_{1},...,f_{s}$ can be output adaptively after $\overrightarrow{\mathbf{m}}_{0}$, $\overrightarrow{\mathbf{m}}_{1}$ are output.
Output: $\mathcal{A}$ output a bit $b^{\prime}$. The adversary $\mathcal{A}$ wins the game if $b^{\prime}=b$. The advantage of $\mathcal{A}$ is the probability that $\mathcal{A}$ wins the game.
We say that the linearly homomorphic signature scheme is $s$-weakly context hiding if the attacker cannot win the privacy game after seeing at most $s$ signatures derived from $\overrightarrow{\mathbf{m}}_{0}$, $\overrightarrow{\mathbf{m}}_{1}$.
3 Prelemimaries
Notation. We use the notations $\mathcal{O}$, $\widetilde{\mathcal{O}}$ and $\omega$ to classify the growth of functions. Let $f,g$ be two function of $n$. $f(n)=\mathcal{O}(g(n))$ if and only if there exits a constant $c$ and an integer $N$ such that $f(n)\leq cg(n)$ for any integer $n>N$. $f(n)=\widetilde{\mathcal{O}}(g(n))$ if and only if $f(n)=\mathcal{O}(g(n)\cdot\log^{c^{\prime}}n)$ for some constant $c^{\prime}$. $f(n)=\omega(g(n))$ if and only if there exists an integer $N$ such that $g(n)\leq cf(n)$ for any constant $c>0$ and any integer $n>N$. Let $n$ denote the security parameter. If $f(n)=\mathcal{O}(n^{c})$ for a constant $c>0$, then we denote $f(n)=\mathrm{poly}(n)$. If for all $c>0$, the function $f(n)$ is $\mathcal{O}(n^{-c})$, then this function is called negligible and denoted as $\mathrm{negl}(n)$. If the probability of an event occurring is $1-\mathrm{negl}(n)$, we say that the event holds with overwhelming probability.
The capital letter $\mathbb{Z}$ denotes the ring of integers, $\mathbb{Z}_{q}$ denotes the ring of integers modulo $q(q\geq 2)$ and $\mathbb{R}^{n}$ denotes the $n$-dimensional Euclidean space. We then let capital letters $A,B,C$, etc. represent matrices, and bold lowercase letters $\mathbf{a},\mathbf{b},\mathbf{c}$ etc. represent vectors. If $A=(\mathbf{a}_{1},...,\mathbf{a}_{n})\in\mathbb{R}^{h\times n}$, the norm of matrix $A$ is defined as $\parallel A\parallel=\max_{1\leq i\leq n}\parallel\mathbf{a}_{i}\parallel$, where $\parallel\mathbf{a}_{i}\parallel$ represents the $l_{2}$-norm of the vector. We further denote $\widetilde{A}=(\widetilde{\mathbf{a}}_{1},...,\widetilde{\mathbf{a}}_{n})$ as the Gram-Schmidt orthogonalization of $A$, that is,
$\widetilde{\mathbf{a}}_{1}=\mathbf{a}_{1},\quad\widetilde{\mathbf{a}}_{i}=\mathbf{a}_{i}-\sum_{j=1}^{i-1}\frac{\langle\mathbf{a}_{i},\widetilde{\mathbf{a}}_{j}\rangle}{\langle\widetilde{\mathbf{a}}_{j},\widetilde{\mathbf{a}}_{j}\rangle}\widetilde{\mathbf{a}}_{j},2\leq i\leq n,$
where $\langle$,$\rangle$ , ⟩ represents the standard inner product in Euclidean space.
For any distribution $\mathcal{D}$, $x\sim\mathcal{D}$ means that $x$ follows the distribution $\mathcal{D}$, and $x\leftarrow\mathcal{D}$ denote that sampling a random value by the distribution. For a set $\mathcal{X}$, let $x\stackrel{{\scriptstyle\$}}{{\leftarrow}}\mathcal{X}$ denote uniformly randomly choosing $x$ form $\mathcal{X}$. For any probabilistic polynomial-time (PPT) algorithm $Alg$, we denote $y\leftarrow Alg(x)$ to mean that the algorithm takes input $x$ and then outputs result $y$.
Definition 6.
(Lattice, [30] Let $\Lambda\subset\mathbb{R}^{n}$ be a non-empty subset. $\Lambda$ is called a lattice if:
(1) it is an additive subgroup of $\mathbb{R}^{n}$;
(2) there exists a positive constant $\lambda=\lambda(\Lambda)>0$ such that
$\mathrm{min}\{\parallel\mathbf{x}\parallel|\mathbf{x}\in\Lambda,\mathbf{x}\neq 0\}=\lambda.$
$\lambda$ is called the minimum distance.
A full-rank $n$-dimensional lattice can also be expressed as a linear combination of a set of basis vectors $B=\{\mathbf{b}_{1},...,\mathbf{b}_{n}\}\subset\mathbb{R}^{n}$:
$\Lambda=\mathcal{L}(B)=\{B\cdot\mathbf{x}=\sum_{i=1}^{n}x_{i}\mathbf{b}_{i}|x=(x_{1},...,x_{n})^{\top}\in\mathbb{Z}^{n}\}.$
We call $\Lambda^{*}$ the dual lattice of $\Lambda$ if
$\Lambda^{*}=\left\{\mathbf{y}\in\mathbb{R}^{n}\mid\langle\mathbf{y},\mathbf{x}\rangle\in\mathbb{Z},\quad\text{for all }\mathbf{x}\in\Lambda\right\}.$
Definition 7.
( $q$-ary lattices, [15]) Let $A\in\mathbb{Z}_{q}^{n\times m}$, $u\in\mathbb{Z}^{n}$. The following two $q$-ary lattices are defined as:
(1)$\Lambda_{q}^{\bot}=\{\mathbf{x}\in\mathbb{Z}^{m}|A\cdot\mathbf{x}\equiv 0(\mathrm{mod}q)\}$
(2)$\Lambda_{q}^{\mathbf{u}}=\{\mathbf{y}\in\mathbb{Z}^{m}|A\cdot\mathbf{y}\equiv\mathbf{u}(\mathrm{mod}q)\}$
The set $\Lambda_{q}^{\mathbf{u}}$ is a coset of $\Lambda_{q}^{\bot}$ since $\Lambda_{q}^{\mathbf{u}}=\Lambda_{q}^{\bot}+\mathbf{t}$ for any $\mathbf{t}$ such that $A\cdot\mathbf{t}\equiv\mathbf{u}(\mathrm{mod}q)$.
Definition 8.
(Short integer solution, [31]) Let $n$, $m$, $q$ be positive integers, with $m=\mathrm{poly(}n)$. Let $A\in\mathbb{Z}_{q}^{n\times m}$ be a uniformly distributed random matrix over $\mathbb{Z}_{q}$, and let $\beta\in\mathbb{R}$ such that $0<\beta<q$. The SIS problem is to find a short integer solution $\mathbf{x}$ satisfying the following condition:
$A\cdot\mathbf{x}\equiv 0(\mathrm{mod}q),\quad\mathrm{and}\quad\mathbf{x}\neq 0,\parallel\mathbf{x}\parallel\leq\beta.$
We denote the SIS problem with parameters $\textbf{SIS}_{q,n,m,\beta}$ or $\textbf{SIS}_{q,\beta}$.
Theorem 1.
(Worst-case to average-case reduction, [32]) In $\mathbb{Z}_{q}^{n\times m}$, $m=poly(n)$, $q\geq\beta\omega(\sqrt{n\log n})$, and $\beta=poly(n)$, if there is a polynomial time algorithm that solves $\textbf{SIS}_{q,\beta}$ with non-negligible probability, then there is a polynomial time algorithm that finds short non-zero vectors, which are only a $\lambda\geq\widetilde{\mathcal{O}}(\sqrt{n})$ factor longer than the shortest vector, in all lattices of dimension $n$.
Definition 9.
(Gaussian distributions, [30]) Let $s$ be a positive real number and $\mathbf{c}\in\mathbb{R}^{n}$ be a vector. The Gaussian function centered at $\mathbf{c}$ with parameter $s$ is defined as: $\rho_{s,\mathbf{c}}(\mathbf{x})=e^{-\frac{\pi}{s^{2}}\parallel\mathbf{x}-\mathbf{c}\parallel^{2}}.$ The discrete Gaussian measure $\mathcal{D}_{\Lambda,s,\mathbf{c}}$ defined on the lattice $\Lambda$ is given by:
$\mathcal{D}_{\Lambda,s,\mathbf{c}}(\mathbf{x})=\frac{\rho_{s,\mathbf{c}}(\mathbf{x})}{\rho_{s,\mathbf{c}}(\Lambda)},$
where $\rho_{s,\mathbf{c}}(\Lambda)=\sum_{\mathbf{x}\in\Lambda}\rho_{s,\mathbf{c}}(x)$.
Definition 10.
(Statistical distance, [30]) Let $M\subset\mathbb{R}^{n}$ be a finite or countable set, and let $X$ and $Y$ be two discrete random variables taking values in $M$. The statistical distance between $X$ and $Y$ is defined as:
$\bigtriangleup(X,Y)=\frac{1}{2}\sum_{a\in M}|P\{X=a\}-P\{Y=a\}|.$
When the statistical distance between two distributions is less than a negligible value, we say that the two distributions are statistically close.
Definition 11.
([33], Min-entropy) For a random variable $X$, its min-entropy is defined as:
$H_{\infty}(X)=-\log\left(\max_{x\in X}\Pr[X=x]\right).$
The average conditional min-entropy of a random variable $X$ conditional on a correlated variable $Y$ is defined as:
$H_{\infty}(X|Y)=-\log\left(\mathbb{E}_{y\in Y}\left\{\max_{x\in X}\Pr[X=x|Y=y]\right\}\right).$
The optimal probability of an unbounded adversary guessing $X$ given the correlated value $Y$ is $2^{-H_{\infty}(X|Y)}$[33].
Lemma 1.
([33]) Let $X$, $Y$ be arbitrarily random variables where the support of $Y$ lies in $\mathcal{Y}$, Then
$H_{\infty}(X|Y)>H_{\infty}(X)-\log(|\mathcal{Y}|).$
Micciancio and Goldwasser [34] proved that a full-rank set $S$ in the lattice $\Lambda$ can be transformed into a basis $T$ such that both have similarly low Gram-Schmidt norms.
Lemma 2.
([34], Lemma 7.1) Let $\Lambda$ be an $n$-dimensional lattice. There is a deterministic polynomial-time algorithm that, given an arbitrary basis of and a full-rank set $S=\{s_{1},...,s_{n}\}$ in $\Lambda$, returns a basis $T$ of $\Lambda$ satisfying
$\parallel\widetilde{T}\parallel\leq\parallel\widetilde{S}\parallel\text{and}\parallel T\parallel\leq\parallel S\parallel\frac{\sqrt{n}}{2}.$
Theorem 2.
[…]
$\parallel\widetilde{T_{A}}\parallel\leq\mathcal{O}(\sqrt{h\log q})\quad\text{and}\quad\parallel T_{A}\parallel\leq\mathcal{O}(h\log q).$
Lemma 3.
(Sampling from discrete Gaussian, [32]) Let $q\geq 2$ , $A\in\mathbb{Z}_{q}^{n\times m}$ and let $T_{A}$ be basis for $\Lambda_{q}^{\bot}(A)$ and $s\geq\widetilde{T_{A}}\cdot\omega(\sqrt{\log m})$. Then for $\mathbf{c}\in\mathbb{R}^{n}$ and $\mathbf{u}\in\mathbb{Z}^{n}$:
  1. 1.
    There is a probabilistic polynomial-time algorithm SampleGaussian$(A,T_{A},s,\mathbf{c})$ that outputs $\mathbf{x}\in\Lambda_{q}^{\bot}(A)$ drawn from a distribution statistically close to $\mathcal{D}_{\Lambda_{q}^{\bot}(A),s,\mathbf{c}}$
  2. 2.
    There is a probabilistic polynomial-time algorithm SamplePre$(A,T_{A},\mathbf{u},s)$ that outputs $\mathbf{x}\in\Lambda_{q}^{\mathbf{u}}(A)$ sampled from a distribution statistically close to $\mathcal{D}_{\Lambda_{q}^{\mathbf{u}}(A),s}$, whenever $\Lambda_{q}^{\mathbf{u}}(A)$ is not empty.
When $s\geq\omega(\sqrt{\log n})$, we denote the Gaussian sampling algorithm over the integer lattice $\mathbb{Z}^{n}$ as $\textbf{SampleDom}(1^{n},s)$. That is, when $\mathbf{x}\stackrel{{\scriptstyle\$}}{{\leftarrow}}\textbf{SampleDom}(1^{n},s)$, $\mathbf{x}$ is statistically close to the distribution $\mathcal{D}_{\mathbb{Z}^{n},s}$. And these preimages have a conditional min-entropy of $\omega(\log n)$[32].
Lemma 4.
([32]) Let $n$ and $h$ be positive integers, and $q$ be a prime, such that $n\geq 2h\lg q$. Then for all but a $2q^{-h}$ fraction of all $A\in\mathbb{Z}_{q}^{h\times n}$ and for any $s\geq\omega(\sqrt{\log n})$, the distribution of the $\mathbf{\alpha}=A\cdot\mathbf{x}(\mathrm{mod}q)$ is statistically close to uniform over $\mathbb{Z}_{q}^{h}$, where $\mathbf{x}\stackrel{{\scriptstyle\$}}{{\leftarrow}}\textbf{SampleDom}(1^{n},s)$.
Lemma 5.
([35]) Let $\Lambda$ be an $n$-dimensional lattice, and $T$ be a basis of the lattice $\Lambda$. If $s\geq\parallel\widetilde{T}\parallel\cdot\omega(\sqrt{\log n})$, then for any $\mathbf{c}\in\mathbb{R}^{n}$, we have:
$\mathrm{Pr}\{\parallel\mathbf{x}-\mathbf{c}\parallel>s\sqrt{n}:\mathbf{x}{\leftarrow}\mathcal{D}_{\Lambda,s,\mathbf{c}}\}\leq\mathrm{negl}(n).$
Definition 12.
( Smoothing parameter[35]) For any $n$-dimensional lattice $\Lambda$ and any given $\epsilon>0$, the smoothing parameter of the lattice is defined as
$\eta_{\epsilon}(\Lambda)=\min\left\{s>0\mid\rho_{\frac{1}{s}}(\Lambda^{*})<1+\epsilon\right\}.$
Lemma 6.
([32]) Let $q\geq 3$, $h$ and $n$ be positive integers satisfying $n\geq 2h\lg q$. Then there exists a negligible function $\epsilon(n)$ such that $\eta_{\epsilon}(\Lambda_{q}^{\bot}(A))<\omega(\sqrt{\log n})$ for all but at most a $q^{-h}$ fraction of $A$ in the $\mathbb{Z}_{q}^{h\times n}$
Lemma 7.
([14]) $\boldsymbol{t}_{i}\in\mathbb{Z}^{m}$ and $\boldsymbol{x}_{i}$ are mutually independent random variables sampled from a Gaussian distribution $D_{\boldsymbol{t}_{i}+\Lambda,\sigma}$ over $\boldsymbol{t}_{i}+\Lambda$ for $i=1,2,\cdots,k$ in which $\Lambda$ is a lattice and $\sigma\in\mathbb{R}$ is a parameter. Let $\boldsymbol{c}=(c_{1},\cdots,c_{k})\in\mathbb{Z}^{k}$ and $g=\mathrm{gcd}(c_{1},\cdots,c_{k})$, $\boldsymbol{t}=\sum_{i=1}^{k}c_{i}\boldsymbol{t}_{i}$. If $\sigma>\|\boldsymbol{c}\|\eta_{\epsilon}(\Lambda)$ for some negligible number $\epsilon$, then $\boldsymbol{z}=\sum_{i=1}^{k}c_{i}\boldsymbol{x}_{i}$ statistically closes to $D_{\boldsymbol{t}+g\Lambda,\|\boldsymbol{c}\|\sigma}$.
Theorem 3.
([14]) Let $\Lambda$ be an $n$-dimensional lattice, $A\in\mathbb{Z}^{s\times k}$, and $T\in\mathbb{Z}^{k\times n}$. $X$ is a random matrix of size $k\times n$ where the rows are independent, and the $i$-th row follows the distribution $\mathcal{D}_{\Lambda+\mathbf{t}_{i},\sigma}$, with $\mathbf{t}_{i}$ being the $i$-th row of the matrix $T$. If $\sigma>\mu(A)\eta_{\epsilon}(\Lambda)$ where $\epsilon$ is some negligible value and $\mu(A)=\sqrt{\mathrm{det}(AA^{\top})}$. Then the random variable $AX$ is statistically close to a distribution parameterized by $\Lambda$, $\sigma$, $A$, $AT$.
This theorem shows that for any matrix $A\in\mathbb{Z}^{s\times k}$ and sufficiently large $\sigma$, the random variable $AX$ is statistically close to a distribution parameterized by $\Lambda,\sigma,A$ ,$AT$. Furthermore, the dependence on the matrix $T$ is only reflected through the quantity $AT$. Therefore, if $X_{1}$ and $X_{2}$ are random variables defined in the same way as $X$, but using matrices $T_{1}$ and $T_{2}$, then when $AT_{1}=AT_{2}$, $AX_{1}$ and $AX_{2}$ are statistically close.
Agrawal et al. [24] noted that the algorithm ExtBasis used in the lattice basis delegation technique by Cash et al.[17] generates a new lattice with a dimension larger than the original one. Therefore, if this lattice basis delegation algorithm is used multiple times, the resulting lattice will have a very large dimension, significantly reducing the efficiency of lattice-based cryptographic schemes designed using this technique. To address this issue, Agrawal et al. developed a new lattice basis delegation technique that maintains a constant dimension during the delegation process. This technique requires the use of a so-called $\mathbb{Z}_{q}$-invertible matrix. As follows, we present the conclusion of Agrawal et al.[24] .
Definition 13.
([24]) (1) We say that a matrix $R$ is $\mathbb{Z}_{q}$-invertible, if this matrix $R$ is invertible in $\mathbb{Z}_{q}^{n\times n}.$ (2) Let $s=\sqrt{h\log q}\cdot\omega(\sqrt{\log n})$. We let $\mathcal{D}_{n\times n}$ denote the distribution on matrices in $\mathbb{Z}^{n\times n}$ defined as $(\mathcal{D}_{(\mathbb{Z}^{n},s)})^{m}$ conditioned on the resulting matrix being $\mathbb{Z}_{q}$-invertible.
Theorem 4.
([24]) There exists a PPT algorithm BasisDel$(A,R,T_{A},s)$. Its inputs are: a matrix $A\in\mathbb{Z}_{q}^{h\times n}$ of rank $h$, a $\mathbb{Z}_{q}$-invertible matrix $R$ sampled from $\mathcal{D}_{n\times n}$, $T_{A}$ is a basis of $\Lambda_{q}^{\bot}(A)$, $s>\parallel\tilde{T}_{A}\parallel n\omega(\log^{2}n)$. The output of this algorithm is a basis $T_{B}$ of the lattice $\Lambda_{q}^{\bot}(B)$, where $B=AR^{-1}$, and $T_{B}<s/\omega(\sqrt{\log n})$.
Inspired by [24], we take a special case of the above algorithm BasisDel$(A,R,T_{A},s)$ that is, replace the input $R$ with an orthogonal matrix $H\in\mathbb{Z}^{n\times n}$. We can obtain a new algorithm NewBasis$(A,H,T_{A})$. Before presenting our algorithm, we first provide a necessary lemma.
Lemma 8.
Let $A\in\mathbb{R}^{n\times n}$ be an $n$-dimensional full-rank matrix, and $H\in\mathbb{R}^{n\times n}$ be an orthogonal matrix. Then,
$\parallel HA\parallel=\parallel A\parallel\quad\text{and}\quad\parallel\widetilde{HA}\parallel=\parallel\widetilde{A}\parallel.$
Proof.
(1) Suppose $A=(\mathbf{a}_{1},...,\mathbf{a}_{n})$, then $HA=(H\mathbf{a}_{1},...,H\mathbf{a}_{n})$. For any $i=1,...,n$ , we have
$\parallel H\mathbf{a}_{i}\parallel^{2}=(H\mathbf{a}_{i})^{T}(H\mathbf{a}_{i})=\mathbf{a}_{i}^{T}H^{T}H\mathbf{a}_{i}=\mathbf{a}_{i}^{T}\mathbf{a}_{i}=\parallel\mathbf{a}_{i}\parallel^{2}.$
Therefore, $\parallel HA\parallel=\parallel A\parallel.$
(2) Let $\widetilde{A}=(\widetilde{\mathbf{a}}_{1},...,\widetilde{\mathbf{a}}_{n}),$ where the vector group $\widetilde{\mathbf{a}}_{1},...,\widetilde{\mathbf{a}}_{n}$ is obtained by Gram-Schmidt orthogonalization of the vector group $\widetilde{\mathbf{a}}_{1},...,\widetilde{\mathbf{a}}_{n}$. Then we have $\widetilde{HA}=(H\widetilde{\mathbf{a}}_{1},...,H\widetilde{\mathbf{a}}_{n})$. The reasons are as follows.
We will prove this by mathematical induction. Given vectors $\mathbf{a}_{1},...,\mathbf{a}_{n}$, according to the definition of Gram-Schmidt orthogonalization:
$\widetilde{\mathbf{a}}_{1}=\mathbf{a}_{1},\quad\widetilde{\mathbf{a}}_{j}=\mathbf{a}_{j}-\sum_{i=1}^{j-1}\frac{\langle\mathbf{a}_{j},\widetilde{\mathbf{a}}_{i}\rangle}{\langle\widetilde{\mathbf{a}}_{i},\widetilde{\mathbf{a}}_{i}\rangle}\widetilde{\mathbf{a}}_{i}\quad j=2,...,n.$
Let $(\widetilde{HA})_{i}$ denote the $i$-th column vector of the matrix $\widetilde{HA}$. Since $\widetilde{\mathbf{a}}_{1}=\mathbf{a}_{1}$, we have $H\widetilde{\mathbf{a}}_{1}=H\mathbf{a}_{1}$. Moreover, the first column vector of $\widetilde{HA}$ is also $H\mathbf{a}_{1}$. Therefore, the first column vector of $\widetilde{HA}$ is identical to that of $H\widetilde{\mathbf{a}}_{1}$.
Assume that for $j=2,...,n-1$, we have $(\widetilde{HA})_{j}=H\widetilde{\mathbf{a}}_{j}$. We can easily prove that $(\widetilde{HA})_{n}=H\widetilde{\mathbf{a}}_{n},$ since
[…]
Therefore, $\widetilde{HA}=(H\widetilde{\mathbf{a}}_{1},...,H\widetilde{\mathbf{a}}_{n})$, and consequently, $\parallel\widetilde{HA}\parallel=\parallel\widetilde{A}\parallel.$
Theorem 5.
Let $q\geq 3$ be odd, $h$ be a positive integer, and let $n:=\lceil 6h\log q\rceil.$ There exists a deterministic polynomial-time algorithm denoted as NewBasis$(A,H,T_{A})$ Its inputs are: a matrix $A\in\mathbb{Z}_{q}^{h\times n}$ of rank $h$, a orthogonal matrix $H\in\mathbb{Z}^{n\times n}$, and $T_{A}$ which is a basis of $\Lambda_{q}^{\bot}(A)$. The output of this algorithm is a basis $T_{B}$ of the lattice $\Lambda_{q}^{\bot}(B)$, where $B=AH^{\top}$, and $\parallel\widetilde{T}_{B}\parallel<\parallel\widetilde{T}_{A}\parallel$, $\parallel T_{B}\parallel<\parallel T_{A}\parallel\sqrt{n}/2$
Proof.
  1. (1)
    According to Theorem 2, run the algorithm $\textbf{TrapGen}(q,h,n)$ to obtain matrix $A\in\mathbb{Z}_{q}^{h\times n}$ and $T_{A}$, where $T_{A}$ is the basis of $\Lambda_{q}^{\bot}(A)$. Take $A$ and $T_{A}$ as the inputs of the algorithm.
  2. (2)
    Select an orthogonal matrix $H$ from $\mathbb{Z}^{n\times n}$. For example, one can choose a diagonal matrix with diagonal entries equal to ±$\pm 1$. It is evident that such a matrix is an orthogonal matrix.
  3. (3)
    We regard $T=HT_{A}$ as a set of linearly independent vectors of the lattice $\Lambda_{q}^{\bot}(B)$, where $B=AH^{\top}$. According to Lemma 2 and Lemma 8, by inputting $T$ and an arbitrary basis of the lattice $\Lambda_{q}^{\bot}(B)$, a basis $T_{B}$ of the lattice $\Lambda_{q}^{\bot}(B)$ can be obtained in polynomial time, satisfying $\parallel\widetilde{T}_{B}\parallel<\parallel\widetilde{HT}_{A}\parallel=\parallel\widetilde{T_{A}}\parallel,\parallel T_{B}\parallel<\parallel HT_{A}\parallel\sqrt{n}/2=\parallel T_{A}\parallel\sqrt{n}/2$.
Remark: for any $q$-ary lattice, a basis can be found in polynomial time without any restrictions on the length of the basis vectors.
At the end of this section, we will present the following two useful lemmas.
Lemma 9.
Let $H$ be an orthogonal matrix over $\mathbb{Z}^{n\times n}$, that is, $HH^{\top}=I_{n}$. If $\mathbf{x}\sim\mathcal{D}_{\mathbb{Z}^{n},\sigma}$, then $H\mathbf{x}\sim\mathcal{D}_{\mathbb{Z}^{n},\sigma}$.
Proof.
(1) Clearly, $H\cdot\mathbb{Z}^{n}=\{H\mathbf{x}|\mathbf{x}\in\mathbb{Z}^{n}\}$ is a lattice. Next, we prove that $\mathbb{Z}^{n}=H\cdot\mathbb{Z}^{n}$.
For any $\mathbf{x}\in\mathbb{Z}^{n}$, we have $\mathbf{x}=H\cdot(H^{\top}\cdot\mathbf{x})$. Since $H^{\top}\cdot\mathbf{x}\in\mathbb{Z}^{n}$, then $\mathbf{x}\in H\cdot\mathbb{Z}^{n}$. Conversely, for any $\mathbf{x}\in H\cdot\mathbb{Z}^{n}$, obviously $\mathbf{x}\in\mathbb{Z}^{n}$. Therefore, $\mathbb{Z}^{n}=H\cdot\mathbb{Z}^{n}$.
(2) Next, we prove that the probability densities of $\mathbf{x}$ and $H\mathbf{x}$ are the same. By the definition of the discrete Gaussian measure:
$\mathcal{D}_{\mathbb{Z}^{n},s}(\mathbf{x})=\frac{\rho_{s}(\mathbf{x})}{\rho_{s}(\mathbb{Z}^{n})}=\frac{e^{-\frac{\pi}{s^{2}}\|\mathbf{x}\|^{2}}}{\rho_{s}(\mathbb{Z}^{n})}=\frac{e^{-\frac{\pi}{s^{2}}\|H\mathbf{x}\|^{2}}}{\rho_{s}(H\cdot\mathbb{Z}^{n})}=\mathcal{D}_{H\cdot\mathbb{Z}^{n},s}(H\mathbf{x}).$
Therefore, $H\mathbf{x}\sim\mathcal{D}_{\mathbb{Z}^{n},s}$. ∎
Lemma 10.
Let $A\in\mathbb{Z}_{q}^{n\times m},s>0,\mathbf{u}\in\mathbb{Z}^{m}$, and $\Lambda=\Lambda_{q}^{\boldsymbol{u}}(A)$. If $\mathbf{x}$ is sampled from $\mathcal{D}_{\mathbb{Z}^{m},s}$ conditioned on $A\mathbf{x}\equiv\boldsymbol{u}\pmod{q}$, then the conditional distribution of $\mathbf{x}$ is $\mathcal{D}_{\Lambda,s}$.
Proof.
According to Definition 9, the discrete Gaussian distribution defined on $\Lambda$ is:
$\mathcal{D}_{\Lambda,s}(x)=\frac{\rho_{s}(\mathbf{x})}{\rho_{s}(\Lambda)},\quad\mathbf{x}\in\Lambda,\text{ where }\rho_{s}(\mathbf{x})=e^{-\frac{\pi}{s^{2}}\|\mathbf{x}\|^{2}}.$
Then, for $\mathbf{x}\sim\mathcal{D}_{\mathbb{Z}^{m},s}$ and $\mathbf{x}\in\Lambda$, it can be regarded as the following conditional distribution:
$\mathcal{D}_{\mathbb{Z}^{m},s}(\mathbf{x}\mid\mathbf{x}\in\Lambda)=\frac{\mathcal{D}_{\mathbb{Z}^{m},s}(\mathbf{x})}{\mathcal{D}_{\mathbb{Z}^{m},s}(\Lambda)}=\frac{\rho_{s}(x)/\rho_{s}(\mathbb{Z}^{m})}{\rho_{s}(\Lambda)/\rho_{s}(\mathbb{Z}^{m})}=\frac{\rho_{s}(\mathbf{x})}{\rho_{s}(\Lambda)}=\mathcal{D}_{\Lambda,s}(\mathbf{x}).$
Therefore, $\mathbf{x}\sim\mathcal{D}_{\Lambda,s}$.
4 New linearly homomorphic signature Scheme
In this section, we will present our linearly homomorphic signature Scheme and prove its correctness.
4.1 Construction
The algorithms TrapGen, NewBasis, and SamplePre are derived from Theorem 2, Theorem 5, and Lemma 3, respectively. Let $\langle f\rangle=(c_{1},...,c_{k})$ denote the linear function $f(\mathbf{m}_{1},...,\mathbf{m}_{k})=c_{1}\mathbf{m}_{1}+...+c_{k}\mathbf{m}_{k}$, where $\mathbf{m}_{i}\in\mathbb{Z}_{2}^{k}$ and $c_{i}\in\{0,1\}$ for all $i$.
[…]
Setup($1^{n}$, k): Input $(n,k)$, where $n$ is the security parameter and $k=\mathcal{O}(n)$ is the maximum size of the data set (the maximum number of linear operations). The parameters are generated as follows:
Let prime $q=\mathrm{poly}(n)$ such that $q\geq(nk)^{3}.$ Let $h=\lfloor\frac{n}{6\log q}\rfloor$, $V=k\sqrt{2n\log q}\cdot\log n$. Output public parameters $pp=(q,h,V)$.
KeyGen($pp$): Input the public parameters $pp$. The algorithm generates a pair of public key and secret key as follows:
  1. (1)
    Compute $(A,T_{A})\longleftarrow\textbf{TrapGen}(q,h,n)$, where $A\in\mathbb{Z}_{q}^{h\times n}$ has rank $h$ and $T_{A}$ is a basis for the lattice $\Lambda_{q}^{\bot}(A)$.
  2. (2)
    Sample $\alpha_{1},\ldots,\alpha_{k}\stackrel{{\scriptstyle\$}}{{\leftarrow}}\mathbb{Z}_{q}^{h}$. We require that $\alpha_{1},\ldots,\alpha_{k}$ are linearly independent. If not, resampling should be carried out, and this process can be easily accomplished within polynomial time.
  3. (3)
    Output the public key $pk=(A,\alpha_{1},\ldots,\alpha_{k})$ and the secret key $sk=T_{A}$.
Sign$(sk,\mathbf{\tau},\mathbf{m}_{i},i)$: Input the secret key $sk=T_{A}$, the tag $\mathbf{\tau}\in\{0,1\}^{n}$, the data message $\mathbf{m}_{i}=(\mathbf{m}_{i1},...,\mathbf{m}_{ik})\in\mathbb{Z}_{2}^{k}$ and the index $i$ indicating that $\mathbf{m}_{i}$ is the $i$-th data message in the data set. the signing algorithm proceeds as follows:
  1. (1)
    Let $H_{\tau}=\mathrm{diag}\{2\tau_{1}-1,2\tau_{2}-1,...,2\tau_{n}-1\}$, then $H_{\tau}$ is an orthogonal matrix.
  2. (2)
    Compute $T_{B^{\tau}}{\leftarrow}\textbf{NewBasis}(A,H_{\tau},T_{A})$, where $T_{B^{\tau}}$ is a basis for $\Lambda_{q}^{\bot}(B^{\tau})$ and $B^{\tau}=AH_{\tau}^{\top}$. Let $\Lambda^{\tau}=\Lambda_{q}^{\bot}(B^{\tau})$.
  3. (3)
    Compute $\mathbf{t}_{i}=\sum_{j=1}^{k}\mathbf{m}_{j}\alpha_{ij}\in\mathbb{Z}_{q}^{h}$
  4. (4)
    Compute $\mathbf{\sigma}_{i}\longleftarrow\textbf{SamplePre}(\Lambda^{\tau},T_{B^{\tau}},\mathbf{t},V)$.
  5. (5)
    Output $\mathbf{\sigma}_{i}$ as signature.
Verify$(pk,\mathbf{\tau},\mathbf{m},\mathbf{\sigma},f)$: Input the public key $pk$, a binary vector $\tau\in\{0,1\}^{n}$, a message $\mathbf{m}\in\mathbb{Z}_{2}^{k}$, a signature $\sigma\in\mathbb{Z}^{n}$, and a linear function $f=(c_{1},\ldots,c_{k})$, where each $c_{i}\in\{0,1\}$. The message $\mathbf{m}$ is defined as $\mathbf{m}=f(\mathbf{m}_{1},\ldots,\mathbf{m}_{k})$. The verification process is as follows:
  1. (1)
    Based on the input $pk=(A,\alpha_{1},\ldots,\alpha_{k})$ and $\mathbf{\tau}$, compute $H_{\tau}=\mathrm{diag}\{2\tau_{1}-1,2\tau_{2}-1,...,2\tau_{n}-1\}$ and $B^{\tau}=AH_{\tau}^{\top}$.
  2. (2)
    Compute $\mathbf{t}_{i}=\sum_{j=1}^{k}\mathbf{m}_{ij}\alpha_{j}$ for $i=1,....k.$
  3. (3)
    Output 1 if all of the following three conditions are satisfied; otherwise, output 0.
    1. 1)
      $\parallel\mathbf{\sigma}\parallel\leq k\cdot V\cdot\sqrt{kn}$;
    2. 2)
      $B^{\tau}\cdot\mathbf{\sigma}(\mathrm{mod}q)=c_{1}\mathbf{t}_{1}+...c_{k}\mathbf{t}_{k}$.
Evaluate$(pk,\mathbf{\tau},f,\overrightarrow{\mathbf{\sigma}})$: Input the public key $pk$, $\mathbf{\tau}\in\{0,1\}^{n}$, a set of signatures $\overrightarrow{\mathbf{\sigma}}=(\sigma_{1},...,\sigma_{k})$, where $\sigma_{i}=\textbf{Sign}(sk,\mathbf{\tau},\mathbf{m},i)$ and the linear function $\langle f\rangle=(c_{1},...,c_{k})$. The evaluation algorithm outputs $\sigma=c_{1}\sigma_{1}+...+c_{k}\sigma_{k}$ as the signature of $f(\mathbf{m}_{1},...,\mathbf{m}_{k})$.
The following proves that the above scheme satisfies correctness.
Theorem 6.
The above linear homomorphic signature scheme satisfies correctness with overwhelming probability.
Proof.
(1) Given $\sigma=\textbf{Sign}(sk,\mathbf{\tau},\mathbf{m}_{i},i)$, let $\pi_{i}$ be the projection function, which can be viewed as a special type of linear function, that is, $\pi_{i}(\mathbf{m}_{1},...,\mathbf{m}_{k})=\mathbf{m}_{i}$. In this case, we check the following two verification conditions:
From Theorem 2 and Theorem 5, it follows that $\parallel\widetilde{T}_{B^{\tau}}\parallel\leq\parallel\widetilde{T}_{A}\parallel\leq\mathcal{O}(\sqrt{h\log q})$ with overwhelming probability. Since $V=k\sqrt{2n\log q}\cdot\log n$, it follows that
$\frac{V}{\parallel\widetilde{T}_{B^{\tau}}\parallel}\geq k\sqrt{\frac{2n}{h}}\cdot\log n\geq\sqrt{\log n},$
which means $V\geq\parallel\widetilde{T}_{B^{\tau}}\parallel\cdot\omega(\sqrt{\log n})$. Therefore, by Lemma 5, we have $\parallel\sigma_{i}\parallel\leq V\cdot\sqrt{kn}<k\cdot V\cdot\sqrt{kn}$ with overwhelming probability.
Since $\sigma_{i}\in\Lambda^{\tau}+\mathbf{t}_{i}$, where $\mathbf{t}_{i}=\sum_{j=1}^{k}\mathbf{m}_{ij}\alpha_{j}$ , thus $B^{\tau}\cdot\sigma_{i}(\mathrm{mod}q)=\mathbf{t}_{i}$.
(2) Given $\sigma_{i}=\textbf{Sign}(sk,\tau,\mathbf{m}_{i},i)$, $(\mathbf{m}_{1},...,\mathbf{m}_{k})\in(\mathbb{Z}_{2}^{k})^{k}$, $f=(c_{1},...,c_{k})$, where $c_{i}\in\{0,1\}$. Let $\overrightarrow{\sigma}=(\sigma_{1},...,\sigma_{k})$. According to the definition of the algorithm $\textbf{Evaluate}(pk,\tau,f,\vec{\sigma})$, $\sigma=c_{1}\sigma_{1}+...+c_{k}\sigma_{k}$ is the signature of $f(\mathbf{m}_{1},...,\mathbf{m}_{k})$. In this case, we will verify that $\sigma$ satisfies the following two verification conditions:
Since $\parallel\sigma_{i}\parallel\leq V\cdot\sqrt{n}$ holds with overwhelming probability, and $|c_{i}|\leq 1$, we have
$\parallel\sigma\parallel=\parallel c_{1}\sigma_{1}+...+c_{k}\sigma_{k}\parallel\leq k\cdot\mathrm{max}_{1\leq i\leq k}\parallel\sigma_{i}\parallel\leq k\cdot V\cdot\sqrt{kn}$
Since $\sigma_{i}=\textbf{Sign}(sk,\tau,\mathbf{m},i)$, we have $B^{\tau}\cdot\sigma_{i}\pmod{q}=\mathbf{t}_{i}$. therefore $B^{\tau}\cdot c_{i}\sigma_{i}\pmod{q}=c_{i}\mathbf{t}_{i}$. Thus,
$B^{\tau}\cdot\sigma\pmod{q}$ $=B^{\tau}\cdot\left(\sum_{i=1}^{k}c_{i}\sigma_{i}\right)\pmod{q}$
$=\sum_{i=1}^{k}B^{\tau}\cdot c_{i}\sigma_{i}\pmod{q}$
$=\sum_{i=1}^{k}c_{i}\mathbf{t}_{i}.$
5 Security and privacy
In the field of provable security, the security of signature algorithms is usually related to hard computational problems. Typically, such proofs describe a reduction from a certain (assumed) hard computational problem to breaking the security properties defined by the signature system. Suppose there exists an adversary $\mathcal{A}$ who can break a signature scheme with a non-negligible advantage $\epsilon$ within polynomial time $t$. We construct a simulator $\mathcal{C}$ to solve a difficult computational problem with an advantage of $\epsilon^{\prime}$ within time $t^{\prime}$, where
$t^{\prime}=t+T,\quad\epsilon^{\prime}=\frac{\epsilon}{L}.$
If $(t^{\prime},\epsilon^{\prime})\approx(t,\epsilon)$, or more specifically, if $T=o(t)$ and $L$ is a constant or relatively small, then we say that this reduction process is a tight reduction, and we say that the signature scheme is tightly secure under the assumption that the computational problem is hard. This requires that the tight security of the signature algorithm be independent of the attacker’s parameters, such as the attacker’s number of queries or its security parameter.
In essence, a tight reduction means that a successful attack algorithm can be transformed into an efficient algorithm for solving difficult computational problems without significantly increasing the running time and or significantly reducing the success probability. Therefore, the tightness of the reduction determines the strength of the security guarantee provided by the security proof: a non-tight reduction provides a weaker security guarantee than a tight reduction. Therefore, it is highly desirable for a signature system to have a tight security reduction.
We will now prove that our linearly homomorphic signature scheme is tightly secure under the EUF-CMA security model in the standard model.
Theorem 7.
Suppose $q=\mathrm{poly}(n)$, $k=\mathcal{O}(n)$, $q\geq(nk)^{3}$, $h=\lfloor\frac{n}{6\log q}\rfloor$, $V=k\sqrt{2n\log q}\cdot\log n$, $\beta=2k\cdot V\cdot\sqrt{kn}$. If the $\textbf{SIS}_{q,\beta}$ problem is hard, then the above linearly homomorphic signature scheme is tightly secure under the EUF-CMA security model.
More specifically, let $\mathcal{A}$ be a polynomial-time adversary that makes at most $q_{s}$ signature queries to the simulator $\mathcal{C}$. If the adversary $\mathcal{A}$ breaks the above signature scheme with advantage $\epsilon$ within time $t$, then the simulator can construct a polynomial-time algorithm $\mathcal{B}$ that within $t+\mathcal{O}(q_{s}T_{\textbf{Sign}}+kT_{\textbf{SampleDom}})$ time outputs a solution to the $\textbf{SIS}_{q,\beta}$ problem with nearly the same advantage $\epsilon-\mathrm{negl}(n)$, where $T_{\textbf{Sign}}$ represents the time spent in generating a signature and $T_{\textbf{SampleDom}}$ represents the time for one execution of the sampling algorithm SampleDom.
Proof.
Assume there exists a polynomial-time adversary $\mathcal{A}$ that breaks the aforementioned signature scheme with advantage $\epsilon$ within time $t$. We will now prove that a simulator $\mathcal{C}$ can be constructed, which utilizes the adversary $\mathcal{A}$’s ability to forge signatures to construct an algorithm $\mathcal{B}$ that can output a solution to the $\textbf{SIS}_{q,\beta}$ problem.
Select an input $A$ uniformly, randomly, and independently from $\mathbb{Z}_{q}^{h\times n}$. It is noted that, with overwhelming probability, matrix $A$ is a row-full-rank matrix. Now the simulator $\mathcal{C}$ generates the public key in the following way:
  1. (1)
    The simulator $\mathcal{C}$ runs the algorithm $\textbf{Setup}(1^{n},k)$ to obtain the public parameter $pp=(q,h,V)$.
  2. (2)
    Sample $k$ vectors using the algorithm $\textbf{SampleDom}(1^{n},s)$, where $s=\frac{V}{\sqrt{k}}\geq\omega(\sqrt{\log n})$:
    $\mathbf{x}_{i}\stackrel{{\scriptstyle\$}}{{\leftarrow}}\textbf{SampleDom}(1^{n},s),\quad\text{for}\quad i=1,...,k,$
    and then compute $\mathbf{\alpha}_{1}=A\cdot\mathbf{x}_{1}(\mathrm{mod}q),...,\mathbf{\alpha}_{k}=A\cdot\mathbf{x}_{k}(\mathrm{mod}q)$.
The simulator $\mathcal{C}$ outputs the public key $pk=(A,\alpha_{1},...,\alpha_{k})$ to adversary $\mathcal{A}$.
After receiving the public key $pk$, the adversary selects a sequence of data sets $\mathbf{m}_{1},...,\mathbf{m}_{q_{s}}$ as signature queries, where $\mathbf{m}_{i}=(\mathbf{m}_{i1},...,\mathbf{m}_{ik})\in(\mathbb{Z}_{2}^{k})^{k}$. For each $i$ ($i=1,...,q_{s}$), the simulator $\mathcal{C}$ choose $\tau_{i}$ uniformly from $\{0,1\}^{n}$ and gives to $\mathcal{A}$ the tag $\tau_{i}$. The steps for the simulator to return the signature queries to the adversary $\mathcal{A}$ are as follows:
  1. (1)
    Take $\tau_{i}\in\{\tau_{1},...,\tau_{q_{s}}\}$, $\mathbf{m}_{i}\in\{\mathbf{m}_{1},...,\mathbf{m}_{q_{s}}\}$.
  2. (2)
    Compute $H_{\tau_{i}}=\mathrm{diag}\{2\tau_{i1}-1,2\tau_{i2}-1,...,2\tau_{in}-1\}$
  3. (3)
    Compute $\mathbf{e}_{ij}=H_{\tau_{i}}\cdot\sum_{l=1}^{k}\mathbf{m}_{ij}^{(l)}\mathbf{x}_{l}$, where $\mathbf{m}_{ij}=(\mathbf{m}_{ij}^{(1)},...,\mathbf{m}_{ij}^{(k)}).$ It is noted that, according to Lemma 7, we have $\sum_{l=1}^{k}\mathbf{m}_{ij}^{(l)}\mathbf{x}_{l}\sim\mathcal{D}_{\mathbb{Z}^{n},V}$.
The simulator $\mathcal{C}$ output the signature $\sigma_{ij}=\mathbf{e}_{ij}$.
Next, we prove that the output of the simulator is the same as the distribution in the real signature scheme (within a negligible statistical distance).
In the real scheme, the public-key matrix $A$ is generated by the algorithm TrapGen, and its distribution is statistically close to the uniform distribution. In the simulation scheme, the public-key matrix $A$ is uniformly and randomly selected, and with overwhelming probability, the matrix $A$ is row- full-rank. Therefore, the two are statistically indistinguishable.
Secondly, the simulation of the partial public keys $\alpha_{1},\alpha_{2},...,\alpha_{k}$ is also indistinguishable from the real scheme. This is because, according to Lemma 4, $\alpha_{i}=A\cdot\mathbf{x}_{i}(\mathrm{mod}q)$ $(i=1,...,k)$ is statistically close to being uniformly distributed.
Finally, we demonstrate that the simulated signature is also statistically indistinguishable from the real signature. Since $\mathbf{x}_{i}\stackrel{{\scriptstyle\$}}{{\leftarrow}}\textbf{SampleDom}(1^{n},s)$ and $s=\frac{V}{\sqrt{k}}\geq\omega(\sqrt{\log n})$. Thus, by Lemma 5, $\parallel\mathbf{x}_{i}\parallel\leq\frac{V}{\sqrt{k}}\sqrt{n}$ holds with overwhelming probability. Therefore,
$\parallel\mathbf{e}_{ij}\parallel=\parallel H_{\tau_{i}}\cdot\sum_{l=1}^{k}\mathbf{m}_{ij}^{(l)}\mathbf{x}_{l}\parallel=\parallel\sum_{l=1}^{k}\mathbf{m}_{ij}^{(l)}\mathbf{x}_{l}\parallel\leq V\cdot\sqrt{kn}\leq k\cdot V\cdot\sqrt{kn}.$
In addition, the adversary $\mathcal{A}$ can verify that:
$AH_{\tau_{i}}^{\top}\cdot\sigma_{ij}$ $=AH_{\tau_{i}}^{\top}\cdot\mathbf{e}_{ij}$ (1)
$=A\cdot H_{\tau_{i}}^{\top}H_{\tau_{i}}\cdot\sum_{l=1}^{k}\mathbf{m}_{ij}^{(l)}\mathbf{x}_{l}=A\cdot\sum_{l=1}^{k}\mathbf{m}_{ij}^{(l)}\mathbf{x}_{l}$
$=\sum_{l=1}^{k}\mathbf{m}_{ij}^{(l)}A\cdot\mathbf{x}_{l}=\sum_{l=1}^{k}\mathbf{m}_{ij}^{(l)}\alpha_{l}=\mathbf{t}_{ij}(\mathrm{mod}q)$
According to Lemma 7 and Lemma 9, $\sigma_{ij}=\mathbf{e}_{ij}$ statistically follows the distribution $\mathcal{D}_{\mathbb{Z}^{n},V}$, but also satisfies the condition $\sigma_{ij}=\mathbf{e}_{ij}\in\Lambda_{q}^{\mathbf{t}_{ij}}(A)$. Therefore, according to Lemma 10, $\sigma_{ij}=\mathbf{e}_{ij}$ statistically follows the distribution $\mathcal{D}_{\Lambda_{q}^{\mathbf{t}_{ij}}(A),V}$. Hence, the simulated signature and the real signature are also statistically indistinguishable.
After the adversary $\mathcal{A}$ receives all the queried signatures, it will generate a valid forgery $(\tau^{*},\mathbf{m}^{*},\sigma^{*},f)$, where $\langle f\rangle=(c_{1},...,c_{k})$ , such that
$\textbf{Verify}(pk,\tau^{*},\mathbf{m}^{*},\sigma^{*},f)=1.$
Whether it is a I-type forgery or II-type forgery, the following equation holds:
$B^{\tau^{*}}\cdot\sigma^{*}(\mathrm{mod}q)=AH_{\tau^{*}}^{\top}\cdot\sigma^{*}=\sum_{i=1}^{k}\mathbf{m}_{i}^{\ast}\alpha_{i}=\sum_{i=1}^{k}\mathbf{m}_{i}^{\ast}A\cdot\mathbf{x}_{i}=A\cdot\mathbf{x}^{\ast}(\mathrm{mod}q),$
where $\mathbf{x}^{\ast}=\sum_{i}^{k}\mathbf{m}_{i}^{\ast}\mathbf{x}_{i}$. By Lemma 5, $\parallel\mathbf{x}_{i}\parallel\leq s\cdot\sqrt{n}\leq\frac{V}{\sqrt{k}}\cdot\sqrt{n}$ holds with overwhelming probability, so $\parallel\mathbf{x}^{\ast}\parallel\leq V\cdot\sqrt{kn}\leq k\cdot V\cdot\sqrt{kn}$. By verification condition (1), $\parallel H_{\tau^{*}}^{\top}\cdot\sigma^{*}\parallel=\parallel\sigma^{*}\parallel\leq k\cdot V\cdot\sqrt{kn}$, thus $\parallel H_{\tau^{*}}^{\top}\cdot\sigma^{*}-\mathbf{x}^{\ast}\parallel\leq 2k\cdot V\cdot\sqrt{kn}$.
Let the process of outputting $H_{\tau^{*}}^{\top}\cdot\sigma^{*}-\mathbf{x}^{\ast}$ be denoted as algorithm $\mathcal{B}$. If $H_{\tau^{*}}^{\top}\cdot\sigma^{*}-\mathbf{x}^{\ast}\neq 0$, then the simulator outputs $H_{\tau^{*}}^{\top}\cdot\sigma^{*}-\mathbf{x}^{\ast}$ as a solution to the $\textbf{SIS}_{q,\beta}$, where $\beta=2k\cdot V\cdot\sqrt{kn}$.
Finally, we discuss the probability of $H_{\tau^{*}}^{\top}\cdot\sigma^{*}=\mathbf{x}^{\ast}$. According to the properties of the sampling algorithm, that is, the preimages have a conditional min-entropy of $\omega(\log n)$. Therefore
$\mathrm{Pr}[H_{\tau^{*}}^{\top}\cdot\sigma^{*}=\mathbf{x}^{\ast}]\leq 2^{-\omega(\log n)}\leq\mathrm{negl}(n).$
Consequently, Algorithm B will output a solution to the $\textbf{SIS}_{q,\beta}$ problem with an advantage of $\epsilon-\mathrm{negl}(n).$
Worst-case to average-case reduction. Based on Theorem 1 from the reference [32], when $q\geq\beta\cdot\omega(\sqrt{n\log n})$ holds, the computational hardness of the $\textbf{SIS}_{q,\beta}$ problem is equivalent to that of approximating the SIVP problem in the worst-case setting with an approximation factor of $\beta\cdot\tilde{O}(\sqrt{n})$. In the Setup algorithm, the requirement $q\geq(nk)^{3}$ serves to ensure that $q$ has a large enough value for Theorem 1 to be applicable.
From the proof of Theorem 7, it can be observed that neither a I-type forgery nor a II-type forgery requires the adversary to specify a target tag $\tau_{g}$ in advance. Therefore, our signature scheme is also tightly secure under the U-ST-SCMA security model.
Theorem 8.
Suppose $q=\mathrm{poly}(n)$, $k=\mathcal{O}(n)$, $q\geq(nk)^{3}$, $h=\lfloor\frac{n}{6\log q}\rfloor$, $V=k\sqrt{2n\log q}\cdot\log n$, $\beta=2k\cdot V\cdot\sqrt{kn}$. If the $\textbf{SIS}_{q,\beta}$ problem is hard, then the above linearly homomorphic signature scheme is tightly secure under the U-ST-SCMA security model.
Finally, we prove that our linearly homomorphic signature scheme is weakly context hiding. We have the following theorem.
Theorem 9.
Assume that $V>k^{s/2}\cdot\omega(\sqrt{1+\log n})$, then the above linearly homomorphic signature scheme is weakly context hiding.
Proof.
The challenger runs KeyGen, generating a public-private key pair $(pk,sk)$, and sends it to the adversary. Upon receiving the public-private key pair, assume the adversary outputs $(\mathbf{m}_{0},\mathbf{m}_{1},f_{1},...,f_{s})$ to the challenger, where $\mathbf{m}_{0}=(\mathbf{m}_{01},...,\mathbf{m}_{0k}),\mathbf{m}_{1}=(\mathbf{m}_{11},...,\mathbf{m}_{1k})$ and satisfies $f_{i}(\mathbf{m}_{0})=f_{i}(\mathbf{m}_{1})$, $i=1,2,...,s.$
For convenience, we treat $\mathbf{m}_{b}$ as a matrix over $\mathbb{Z}_{p}^{k\times n}$, where the $j$-th row corresponds to the $j$-th message in the data set $\mathbf{m}_{b}$. We denote $<f_{i}>=(c_{i1},...,c_{ik})$ and construct the $s\times k$ matrix $G$, where the $i$-th row is $<f_{i}>$.Therefore, we have
$G\mathbf{m}_{0}=G\mathbf{m}_{1}.$
In response, the challenger randomly selects a tag $\tau\in\{0,1\}^{n}$ and a random bit $b\in\{0,1\}$. We define $\sigma_{j}^{(b)}$ as the signature of the $j$-th message in the data set $\mathbf{m}_{b}$. We then construct a $k\times n$ matrix $\Sigma^{(b)}$, where the $j$-th row is $(\sigma_{j}^{(b)})^{\top}$. For each $i=1,2,...,s$, let $d_{i}^{(b)}=\textbf{Evaluate}(pk,\tau,f_{i},(\sigma_{1}^{(b)},...,\sigma_{k}^{(b)}))$. Then, we construct an $s\times n$ matrix $D^{(b)}$, where the $j$-th row is $(d_{j}^{(b)})^{\top}$. According to the definition of Evaluate, we have
$D^{(b)}=G\Sigma^{(b)}.$
For $b=0,1$, $\sigma_{i}^{(b)}$ is drawn from $\mathcal{D}_{\Lambda^{\tau}+\mathbf{t}_{i}^{(b)},V}$, where $\mathbf{t}_{i}^{(b)}=\sum_{j=1}^{k}\mathbf{m}_{bi}^{(j)}\alpha_{j}.$ By the definition of the signing algorithm, the signatures are independent of each other. Therefore, when $V$ is sufficiently large $(>\mu(G)\eta_{\varepsilon}(\Lambda^{\tau})$, by Theorem 3, the distributions of $D^{(0)}$ and $D^{(1)}$ are statistically close. Therefore, any polynomial-time adversary cannot distinguish between $D^{(0)}$ and $D^{(1)}$ with non-negligible probability and further guess the value of $b$. Now we only need to prove that
$V>k^{s/2}\cdot\omega(\sqrt{1+\log n})\geq\mu(G)\eta_{\varepsilon}(\Lambda^{\tau}).$
First, since the elements of the matrix $G$ are from the set $\{0,1\}$, then by the Hadamard inequality:
$\mu(G)=\sqrt{det(GG^{\top})}\leq\prod_{i=1}^{s}\parallel\mathbf{g}_{i}\parallel\leq k^{s/2}<V,$
where $\mathbf{g}_{i}$ represents the $i$-th row of the matrix $G$.
Secondly, According to Lemma 6, there exists a negligible $\epsilon$ such that
$\eta_{\epsilon}(\Lambda^{\tau})\leq\omega(\sqrt{1+\log n})<V$
holds with overwhelming probability. Therefore,
$V>k^{s/2}\cdot\omega(\sqrt{1+\log n})\geq\mu(G)\eta_{\varepsilon}(\Lambda^{\tau}).$
6 Conclusion
At the ASIACRYPT 2016 conference, Boyen and Li [20] posed an open question: how to construct a tightly secure short homomorphic signature scheme based on lattices. In 2024, Chen [21] constructed the first lattice-based short and tightly secure linear homomorphic signature scheme. However, Chen’s scheme does not achieve tight security under the standard EUF-CMA model. To achieve tight security, they defined a weaker security model, U-ST-SCMA, and proved tight security under this model.
In contrast, this paper addresses this issue, and our scheme can be proven to achieve tight security under the EUF-CMA . Furthermore, our scheme offers significant improvements over [21] in terms of public key size, signature size, and private key size (see Table 1). Therefore, among existing lattice-based linearly homomorphic signature schemes with tight security, our scheme is the most efficient.
However, several unresolved issues remain, such as how to construct an efficient and tightly secure polynomial function homomorphic signature or fully homomorphic signature based on lattices in the standard model. Additionally, another intriguing question is whether it is possible to construct a secure homomorphic signature over a semigroup, which is left in [8].