### v5

Abstract
Boyen and Li posed an open problem in their ASIACRYPT 2016 conference paper: How to construct a tightly secure homomorphic signature scheme under the Short Integer Solution (SIS) hardness assumption in the standard model. This work provides the first complete resolution of this problem under the same assumption.
keywords:
lattice, linearly homomorphic signature, tightly secure, standard model, SIS problem
1 Introduction
Homomorphic signatures refer to a type of signature that allows any entity to perform homomorphic operations on authenticated data, enabling the generation of new data. The entity can then obtain a valid signature for the new data without using the signing private key. Since homomorphic signatures were put forward, they have attracted more and more attention. The unique properties of homomorphic signatures provide a vast theoretical research space and significant scientific research value, playing an important role in many practical applications. For example, they can offer effective solutions in fields such as network coding[2, 3, 4, 35, 36] and cloud computing[5, 6, 37, 38, 39].
The concept of homomorphic signature schemes was first introduced by R. Rivest [7] in a presentation at Cambridge University in 2000. Shortly thereafter, in 2002, Johnson et al. [8] provided the formal definition and security analysis of homomorphic signatures. Since then, various types of homomorphic signature schemes have emerged. When classifying homomorphic signature schemes based on the homomorphic operation functions, they can be categorized into linear homomorphic signatures, polynomial function homomorphic signatures, and fully homomorphic signatures.
In 2007, Zhao et al. proposed the first linearly homomorphic signature scheme [2], which allows arbitrary linear combinations of signed data, facilitating the verification of message integrity and effectively preventing contamination attacks on applications based on network coding. In 2009, Boneh et al. [9] provided the formal definition and security model for linearly homomorphic signature schemes and constructed a linearly homomorphic signature scheme for packet authentication in network coding protocols based on bilinear groups. Subsequently, further improvements were made in terms of efficiency, security, and privacy protection, leading to the proposal of numerous efficient and practical linearly homomorphic signature schemes. Early research primarily based linearly homomorphic signatures on number-theoretic assumptions, such as the discrete logarithm problem or the RSA problem [9, 10, 11, 12, 13, 40, 41, 42]. These schemes are efficient but must be designed over $\mathbb{F}_{p}$ (where $p$ is a large prime) to ensure the difficulty of the RSA or discrete logarithm problem, and they cannot resist quantum computing attacks.
To resist quantum attacks, Boneh and Freeman[14] first proposed a linearly homomorphic signature scheme over $\mathbb{F}_{2}$ (the binary field) at the PKC 2011 conference, with its security based on the $k$-SIS problem ($k$-small integer solution problem). Subsequently, at the EUROCRYPT 2011 conference, Boneh and Freeman[15] introduced another linearly homomorphic signature scheme, whose security relies on the hardness of the SIS problem, and proposed a homomorphic signature scheme for polynomial functions, with security based on ideal lattices. In 2013, Wang et al.[16] improved the first linearly homomorphic signature scheme based on lattices, resulting in a new linearly homomorphic signature whose security is directly based on the standard SIS problem. All these schemes[14, 15, 16] have been proven secure in the random oracle model.
However, in reality, the ideal random oracle model is difficult to achieve. Chen et al. combined bonsai tree techniques[17] with the intersection method of two integer lattices proposed in [14], and for the first time constructed a lattice-based linearly homomorphic signature scheme over small fields in the standard model[18]. Building on Chen’s work, Lin et al. constructed two lattice-based linearly homomorphic signature schemes in the standard model by utilizing full-rank differential hash functions and linearly homomorphic chameleon hash functions[19].
In the security reduction of a cryptographic system, if the loss in the adversary’s advantage is only a constant factor, the system is said to be tightly secure; if the security loss is a logarithmic factor of the security parameter, it is referred to as almost tightly secure(as discussed in Section 4). Tight security is a highly desirable property in the design of cryptographic schemes, as it provides stronger practical security guarantees. Regarding this security property, Boyen and Li [20] posed an open problem at the 2016 Asiacrypt conference (ASIACRYPT 2016): How to construct a short, homomorphic signature scheme with tight or almost tight security in the standard model, based on the hardness of the Short Integer Solution (SIS) problem. In 2024, Chen et al. [21] designed the first lattice-based almost tightly secure linearly homomorphic signature scheme by building upon the scheme in literature [18], incorporating pseudorandom function techniques[20] and the key-homomorphic algorithm Eval [22, 23]. However, the scheme fails to achieve almost tight security under the existential unforgeability against chosen-message attacks (EUF-CMA) security model(see Definition 3 for details). Under the standard model, constructing a homomorphic signature scheme with tight security based on the SIS problem remains an open problem, particularly in achieving tight security under the standard EUF-CMA security model.
1.1 Our Contribution
In this paper, we address the above-mentioned open problem. More specifically, we propose the first lattice-based short linear homomorphic signature scheme with tight security under the standard model.
Secondly, among the lattice-based linearly homomorphic schemes that are tight or nearly tight under the standard model (only seen in [21]), our scheme demonstrates sufficient advantages in both efficiency and security under the same parameter settings. Although [21] generates short signatures, it exhibits a complex key generation procedure, requiring the random selection of $\mathcal{O}(s+5)$ matrices and $\mathcal{O}(k)$ vectors for the public key, while the private key consists of one matrix and one vector. Furthermore, it employs the Bonsai Tree technique proposed by Cash et al.[17], which results in the final signature dimension being doubled. Compared with [21] under the same parameters, the linearly homomorphic signature scheme we construct requires only $\mathcal{O}(1)$ randomly selected matrices and $\mathcal{O}(k)$ vectors for the public key. The private key, in turn, consists of only one matrix. Moreover, the size of our signature is half the size of that in [21]. In terms of security, [20] demonstrates that its scheme achieves almost tight security under the Unforgeability under selective-tag static chosen-message attack (U-ST-SCMA) model (see Definition 4 for details). In fact, compared with EUF-CMA, this security model is weaker because the attacker is non-adaptive. Our scheme, however, not only satisfies EUF-CMA security but also exhibits tight security. A detailed comparison is presented in Table 1
Table 1: Comparison of the [21]
Schemes Public key Signature Private key Security model Tightness
[21] $\mathcal{O}(s+5)\left\|\mathbb{Z}_{q}^{h\times n}\right\|+\mathcal{O}(k)\left\|\mathbb{Z}_{q}^{h}\right\|$ $\left\|\mathbb{Z}^{2n}\right\|$ $\left\|\mathbb{Z}^{n\times n}\right\|+\left\|\mathbb{Z}_{2}^{s}\right\|$ U-ST-SCMA almost tight
Ours $\mathcal{O}(1)\left\|\mathbb{Z}_{q}^{h\times n}\right\|+\mathcal{O}(k)\left\|\mathbb{Z}_{q}^{h}\right\|$ $\left\|\mathbb{Z}^{n}\right\|$ $\left\|\mathbb{Z}^{n\times n}\right\|$ EUF-CMA tight
1.2 Our Techniques
In 2024, Chen et al. [21] proposed an almost-tight linearly homomorphic signature scheme in the standard model, building upon the approach of Boyen et al. [20]. Their construction utilizes a pseudorandom function (PRF) [20] and a key-homomorphic evaluation algorithm Eval developed in [22, 23]. However, their scheme does not achieve an almost-tight reduction under the EUF-CMA security model. Therefore, to construct a tightly secure scheme under the EUF-CMA model, new technical approaches must be introduced.
In this paper, we employ the trapdoor generation algorithm Trapdoor $\textbf{TrapGen}(q,h,n)$ to construct the scheme. This algorithm generates a matrix $A\in\mathbb{Z}_{q}^{h\times n}$ with an approximately uniform distribution and its corresponding trapdoor $T_{A}\in\mathbb{Z}^{n\times n}$, where $A$ serves as part of the public key while T functions as the private key. To reduce the signature length of the scheme by half compared to that in reference [20], we further design a novel lattice basis delegation algorithm, $\textbf{NewBasis}(A,H,T_{A})$.
Our design motivation is as follows: The signature generation in [18, 21] primarily relies on the lattice basis delegation technique ExtBasis proposed by Cash et al. [17]. This technique requires the introduction of additional public matrices to generate new lattices, which are then incorporated as part of the public key. However, this approach not only results in the dimension of the new lattice being larger than that of the original lattice but also causes the dimension of signatures generated via trapdoor sampling on the new lattice to increase accordingly. In 2010, Agrawal et al. identified this limitation of ExtBasis and proposed an improved lattice basis delegation algorithm, $\textbf{BasisDel}(A,R,T_{A},s)$[24], which maintains a constant dimension during the basis delegation process. In this algorithm, $A\in\mathbb{Z}_{q}^{h\times n}$ of rank $h$, $R\in\mathbb{Z}_{q}^{n\times n}$ is an invertible matrix, $T_{A}$ is the basis of $\Lambda_{q}^{\bot}(A)$, and $s$ is a positive real number. The algorithm outputs a new trapdoor $T_{B}$, where $B=AR^{-1}$.
Theoretically, directly employing BasisDel could significantly reduce the signature length. However, in the security reduction of our proposed scheme, we observed that the norm of the matrix $R^{-1}$ cannot be bounded sufficiently small, thereby preventing the construction of a short vector solution for solving the SIS problem. Furthermore, our construction requires that $R^{-1}$ does not alter the norm of vectors or matrices upon operation. Consequently, we replaced $R$ with an orthogonal matrix $H$ and introduced specific modifications to the BasisDel algorithm, ultimately proposing the NewBasis algorithm.
Another key technique proposed in this paper is the decomposition algorithm, termed $\textbf{Decompose}(\mathbf{m})$. This algorithm decomposes the majority of vectors $\mathbf{m}\in\mathbb{F}_{2}^{2n}$ into a pair of vectors $(\mathbf{u},\mathbf{v})$ with specific Hamming weights, where the Hamming weight of $\mathbf{u}$ is fixed at $n-1$, and that of $\mathbf{v}$ is either $n$ or $n+1$, such that $\mathbf{x}=\mathbf{u}+\mathbf{v}$. The pivotal role of this algorithm lies in the security proof, ensuring that the simulated signatures are statistically indistinguishable from the genuine signatures. This guarantees that the adversary cannot differentiate between the simulated environment and the real environment during forgery attacks. Given that the related analysis involves intricate statistical properties, the detailed discussion will be presented in Theorem 10 (Security Proof).
Based on the aforementioned techniques, our signature scheme first generates a matrix $A$ and its trapdoor $T_{A}$ (serving as the private key) via the trapdoor generation algorithm $\mathsf{TrapGen}(q,h,n)$. Simultaneously, $k$ (an even number) linearly independent vectors $\{\boldsymbol{\alpha}_{1},\ldots,\boldsymbol{\alpha}_{k}\}$ are selected from $\mathbb{Z}_{q}^{h}$ to constitute the public key jointly with $A$.
For a message $m_{i}\in\mathbb{F}_{2}^{k}$ to be signed and its corresponding tag $\tau\in\{0,1\}^{n}$, the tag $\tau$ is first mapped to a diagonal matrix $H_{\tau}=\mathrm{diag}\{2\tau_{1}-1,\ldots,2\tau_{n}-1\}$ (note that $H_{\tau}$ is evidently an orthogonal matrix). The algorithm $\mathsf{NewBasis}(A,H_{\tau},T_{A})$ is then employed to generate a new trapdoor $T_{B^{\tau}}$, where $B^{\tau}=AH_{\tau}^{\top}$.
Subsequently, the message $m_{i}$ is decomposed into a pair of vectors $(\mathbf{u}_{i},\mathbf{v}_{i})$ via the algorithm $\mathsf{Decompose}(m_{i})$. If neither $\mathbf{u}_{i}$ nor $\mathbf{v}_{i}$ is a zero vector, the linear combinations $t(\mathbf{u}_{i})=\sum_{j=1}^{k}u_{ij}\boldsymbol{\alpha}_{j}$ and $t(\mathbf{v}_{i})=\sum_{j=1}^{k}v_{ij}\boldsymbol{\alpha}_{j}$ are computed. The preimage sampling algorithm $\mathsf{SamplePre}(B^{\tau},\mathbf{T}_{B^{\tau}},t(\mathbf{u}_{i}),V)$ and $\mathsf{SamplePre}(B^{\tau},T_{B^{\tau}},t(\mathbf{v}_{i}),V)$ are then invoked to generate the signature components $\boldsymbol{\sigma}(\mathbf{u}_{i})$ and $\boldsymbol{\sigma}(\mathbf{v}_{i})$, respectively. Here, $V$ denotes the Gaussian parameter controlling the sampling distribution.
Finally, the signature is output as $\boldsymbol{\sigma}_{i}=\boldsymbol{\sigma}(\mathbf{u}_{i})+\boldsymbol{\sigma}(\mathbf{v}_{i})$.
2 Preliminaries
2.1 Notation
We use the notations $\mathcal{O}$, $\widetilde{\mathcal{O}}$, and $\omega$ to characterize the growth of functions. Let $f$ and $g$ be two functions of $n$. We say that $f(n)=\mathcal{O}(g(n))$ if and only if there exist a constant $c$ and an integer $N$ such that $f(n)\leq cg(n)$ for all integers $n>N$. Similarly, $f(n)=\widetilde{\mathcal{O}}(g(n))$ holds if and only if $f(n)=\mathcal{O}(g(n)\cdot\log^{c^{\prime}}n)$ for some constant $c^{\prime}$. $f(n)=\omega(g(n))$ if and only if there exists an integer $N$ such that $g(n)\leq cf(n)$ for any constant $c>0$ and any integer $n>N$. Let $n$ denote the security parameter. If $f(n)=\mathcal{O}(n^{c})$ for a constant $c>0$, then we denote $f(n)=\mathrm{poly}(n)$. If for all $c>0$, the function $f(n)$ is $\mathcal{O}(n^{-c})$, then this function is called negligible and denoted as $\mathrm{negl}(n)$. If the probability of an event occurring is $1-\mathrm{negl}(n)$, we say that the event holds with overwhelming probability.
The capital letter $\mathbb{Z}$ denotes the ring of integers, $\mathbb{Z}_{q}$ denotes the ring of integers modulo $q(q\geq 2)$ and $\mathbb{R}^{n}$ denotes the $n$-dimensional Euclidean space. We then let capital letters $A,B,C$, etc. represent matrices, and bold lowercase letters $\mathbf{a},\mathbf{b},\mathbf{c}$ etc. represent vectors. If $A=(\mathbf{a}_{1},...,\mathbf{a}_{n})\in\mathbb{R}^{h\times n}$, the norm of the matrix $A$ is defined as $\parallel A\parallel=\max_{1\leq i\leq n}\parallel\mathbf{a}_{i}\parallel$, where $\parallel\mathbf{a}_{i}\parallel$ represents the $l_{2}$ norm of the vector. We further denote $\widetilde{A}=(\widetilde{\mathbf{a}}_{1},...,\widetilde{\mathbf{a}}_{n})$ as the Gram-Schmidt orthogonalization of $A$, that is,
$\widetilde{\mathbf{a}}_{1}=\mathbf{a}_{1},\quad\widetilde{\mathbf{a}}_{i}=\mathbf{a}_{i}-\sum_{j=1}^{i-1}\frac{\langle\mathbf{a}_{i},\widetilde{\mathbf{a}}_{j}\rangle}{\langle\widetilde{\mathbf{a}}_{j},\widetilde{\mathbf{a}}_{j}\rangle}\widetilde{\mathbf{a}}_{j},2\leq i\leq n,$
where $\langle$,$\rangle$ , ⟩ represents the standard inner product in Euclidean space.
For any distribution $\mathcal{D}$, $x\sim\mathcal{D}$ means that $x$ follows the distribution $\mathcal{D}$, and $x\leftarrow\mathcal{D}$ means that the sampling of a random value by the distribution. For a set $\mathcal{X}$, let $x\stackrel{{\scriptstyle\$}}{{\leftarrow}}\mathcal{X}$ denote uniformly randomly choosing $x$ form $\mathcal{X}$. For any probabilistic polynomial-time (PPT) algorithm $Alg$, we write $y\leftarrow Alg(x)$ to denote that the algorithm takes input $x$ and then outputs the result $y$.
2.2 Linearly homomorphic signature: definition and security model
Before introducing the linearly homomorphic signature, we define a multi-data homomorphic signature [21, 26, 27]. Different from the classical definition of digital signatures, this homomorphic signature consists of five polynomial-time algorithms (Setup, KeyGen, Sign, Verify, Evaluate), allowing the signer to sign multiple different data sets (each data set corresponds to a tag), and the size of each data set can be different. It is also assumed that the verifier knows the tag of the data set over which they want to verify the computation. The definition of the multi-data homomorphic signature is as follows.
Definition 1.
(Multi-data homomorphic signature, [21, 26, 27]) A multi-data homomorphic signature scheme is a tuple of probabilistic, polynomial-time algorithms (Setup, KeyGen, Sign, Verify, Evaluate) as follows:
  1. $\bullet$
    Setup$(1^{n},k)$. The algorithm takes as input a security parameter $n$ and and a maximum data set size $k$, and outputs a public parameter $pp$. The public parameter $pp$ defines a message space $\mathcal{M}$, a signature space $\Sigma$, and a set $\mathcal{F}$ of admissible functions $f:\mathcal{M}^{k}\longrightarrow\mathcal{M}.$
  2. $\bullet$
    KeyGen$(pp)$. Taking the public parameter $pp$ as input, this algorithm outputs a pair of public key and private key $(pk,sk)$
  3. $\bullet$
    Sign$(sk,\tau,\mathbf{m}_{i},i)$. Takes a secret key $sk$, a tag $\tau\in\{0,1\}^{n}$, a message $\mathbf{m}_{i}\in\mathcal{M}$ and index $i\in\{1,...,k\}$, and output a signature $\sigma_{i}\in\Sigma$.
  4. $\bullet$
    Verify$(pk,\tau,\mathbf{m},\sigma,f)$. Takes a public key $pk$, a tag $\tau\in\{0,1\}^{n}$, a message $\mathbf{m}\in\mathcal{M}$ a signature $\sigma\in\Sigma$, and a function $f\in\mathcal{F}$, and outputs either 0 (reject) or 1 (accept).
  5. $\bullet$
    Evaluate$(pk,\tau,f,\vec{\sigma})$. Takes a public key $pk$, a tag $\tau\in\{0,1\}^{n}$, a function $f\in\mathcal{F}$, and a tuple of signatures $\vec{\sigma}\in\Sigma^{k}$, and outputs a signature $\sigma^{\prime}\in\Sigma$.
Let $\pi_{i}:\mathcal{M}^{k}\longrightarrow\mathcal{M}$ be the function $\pi_{i}(\mathbf{m}_{1},...,\mathbf{m}_{k})=\mathbf{m}_{i}$ that projects onto the $i$-th component. We require that $\pi_{1},...,\pi_{k}\in\mathcal{F}$ for all $pk$ output by Setup$(1^{n},k)$.
For correctness, we require that for each $(pk,sk)$ output by Setup$(1^{n},k)$, it holds that:
  1. (1)
    For all tags $\tau\in\{0,1\}^{n}$, all $\mathbf{m}\in\mathcal{M}$, and all $i\in\{1,...,k\}$, if $\sigma\longleftarrow\textbf{Sign}(sk,\tau,\mathbf{m},i)$, then
    $\textbf{Verify}(pk,\tau,\mathbf{m},\sigma,\pi_{i})=1.$
  2. (2)
    For all tags $\tau\in\{0,1\}^{n}$, all $\overrightarrow{\mathbf{m}}=(\mathbf{m}_{1},...,\mathbf{m}_{k})\in\mathcal{M}^{k}$, and all $f\in\mathcal{F}$, if $\sigma_{i}\longleftarrow\textbf{Sign}(sk,\tau,\mathbf{m},i)$, for   $i\in\{1,...,k\}$, then
    $\textbf{Verify}(pk,\tau,f(\overrightarrow{\mathbf{m}}),\textbf{Evaluate}(pk,\tau,f,(\sigma_{1},...,\sigma_{k})))=1.$
We say that a signature scheme as above is multi-data $\mathcal{F}$-homomorphic.
Remark 1: Typically, we require the message dataset $\mathbf{m}_{1},...,\mathbf{m}_{k}$ to be a set of linearly independent vectors.
The definition of linearly homomorphic signatures is given below.
Definition 2.
([21]) In the above definition of a multi-data homomorphic signature scheme, if the message space $\mathcal{M}$ consists of $n$-dimensional vectors over a ring $\mathbf{R}$, and the set of admissible functions $\mathcal{F}$ consists of $\mathbf{R}$-linear function from $(\mathbf{R^{n}})^{k}$ to $\mathbf{R}$, we refer to it as the multi-data linearly homomorphic signature scheme. A linear function $f(\mathbf{v}_{1},...,\mathbf{v}_{k})=\sum_{i=1}^{i=k}c_{i}\mathbf{v}_{i}$ is represented as the vector $(c_{1},...,c_{k})\in\mathbf{R}^{k}$.
Referring to [15, 19], the security model of linearly homomorphic signature is given below. (Note: The signature scheme definition provided in [15, 19] is essentially no different from Definition 2)
Definition 3.
(EUF-CMA, [15, 19]) A linearly homomorphic signature $\mathcal{LS}$=(Setup, Sign, Verify, Evaluate) satisfies existential unforgeability under chosen-message attacks (EUF-CMA) if for all $k$ the advantage of any probabilistic, polynomial-time adversary $\mathcal{A}$ in the following game is negligible in the security parameter $n$:
[…]
Queries: Proceeding adaptively, $\mathcal{A}$ specifies a sequence of data sets $\overrightarrow{\mathbf{m}}_{i}\in\mathcal{M}^{k}$. For each $i$, the challenger choose $\tau_{i}$ uniformly from $\{0,1\}^{n}$ and gives to $\mathcal{A}$ the tag $\tau_{i}$ and the signature $\sigma_{i,j}=\textbf{Sign}(sk,\tau_{i},\mathbf{m}_{ij},j)$ for $j=1,...,k$.
Output: $\mathcal{A}$ output a tag $\tau^{*}\in\{0,1\}^{n}$, a message $\mathbf{m}^{*}\in\mathcal{M}$, a linear function $f\in\mathcal{F}$, and a signature $\sigma^{*}\in\Sigma$.
The adversary wins if $\textbf{Verify}(pk,\tau^{*},\mathbf{m}^{*},\sigma^{*},f)=1$ and either
  1. (1)
    $\tau^{*}\neq\tau_{i}$ for all $i$ (I-type forgery), or
  2. (2)
    $\tau^{*}=\tau_{i}$ for some $i$ but $\mathbf{m}^{*}\neq f(\overrightarrow{\mathbf{m}}_{i})$ (II-type forgery).
The advantage of $\mathcal{A}$ is the probability that $\mathcal{A}$ wins the security game.
In reference [19], the authors also defined a weaker security model, namely existential unforgeability against non-adaptive chosen-message attacks (EUF-naCMA). This model requires the adversary A to submit the set of messages to be signed to the challenger before the attack begins. Due to space limitations and the fact that this model falls outside the scope of this paper, its detailed definition will not be elaborated here. Readers may refer to [19] for further details.
In [21], the purpose of Chen is to construct a almost tightly secure linearly homomorphic signature Scheme. However, in practice, it cannot satisfy almost tight security under EUF-CMA. Therefore, inspired by [28, 29], they proposed unforgeability under the selective-tag static chosen-message-attack (U-ST-SCMA), and achieved tight security under this security model.
Definition 4.
(U-ST-SCMA, [21]) For a linearly homomorphic signature scheme $\mathcal{LS}$=(Setup, Sign, Verify, Evaluate), we consider the following game:
Initialization: (select tag and data phase): Let $n_{s}$ be the number of the data sets . The challenger uniformly samples $(\tau_{i}\stackrel{{\scriptstyle\$}}{{\leftarrow}}\{0,1\}^{n})_{i\in\{1,...,n_{s}\}}$ and sends them to the adversary $\mathcal{A}$. Then, $\mathcal{A}$ specifies a sequence of data sets $(\overrightarrow{\mathbf{m}}_{i}=(\mathbf{m}_{i1},...,\mathbf{m}_{ik})\in\mathcal{M}^{k})_{i\in\{1,...,n_{s}\}}$ and a target tag $\tau_{g}\in\{\tau_{1},...,\tau_{n_{s}}\}$, and returns them.
[…]
Queries: The challenger computes $(\sigma_{ij}{\leftarrow}\textbf{Sign}(sk,\tau_{i},\mathbf{m}_{ij},j))_{i\in\{1,...,n_{s}\},j\in\{1,...,k\}}$, and gives the signatures $(\sigma_{ij})_{i\in\{1,...,n_{s}\},j\in\{1,...,k\}}$ to $\mathcal{A}$.
Output: $\mathcal{A}$ outputs a tag $\tau^{*}\in\{0,1\}^{n}$, a message $\mathbf{m}^{*}\in\mathcal{M}$, a linear function $f\in\mathcal{F}$, and a signature $\sigma^{*}\in\Sigma$. Then The adversary wins if $\textbf{Verify}(pk,\tau^{*},\mathbf{m}^{*},\sigma^{*},f)=1$ and either
  1. (1)
    $\tau^{*}\neq\tau_{i}$ for all $i$ (I-type forgery), or
  2. (2)
    $\tau^{*}=\tau_{g}$ but $\mathbf{m}^{*}\neq f(\overrightarrow{\mathbf{m}}_{g})$ (II-type forgery).
Compared to EUF-CMA, U-ST-SCMA is a weaker security form, as this scheme requires the pre-disclosure of the messages to be signed and the target tags to be forged before providing the public key to the adversary.
In homomorphic signatures, privacy also needs to be considered. Below, we provide the definition of weakly context hiding[15].
Definition 5.
(Weakly context hiding, [15]) A linearly homomorphic signature scheme $\mathcal{LS}$=(Setup, Sign, Verify, Evaluate) is weakly context hiding if for all $k$, the advantage of any probabilistic, polynomial-time adversary $\mathcal{A}$ in the following game is negligible in the security parameter $n$:
[…]
Challenge: $\mathcal{A}$ outputs $(\overrightarrow{\mathbf{m}}_{0},\overrightarrow{\mathbf{m}}_{1},f_{1},...,f_{s})$ with $\overrightarrow{\mathbf{m}}_{0},\overrightarrow{\mathbf{m}}_{1}\in\mathcal{M}^{k}$. The function $f_{1},...,f_{s}\in\mathcal{F}$ and satisfy
$f_{i}(\overrightarrow{\mathbf{m}}_{0})=f_{i}(\overrightarrow{\mathbf{m}}_{1})\quad\text{for all}\quad i=1,...,s.$
In response, the challenger generates a random bit $b\in\{0,1\}$ and a random tag $\tau\in\{0,1\}^{n}$. It signs the message in $\overrightarrow{\mathbf{m}}_{b}$ using the tag $\tau$ to obtain a vector $\overrightarrow{\sigma}$ of $k$ signatures. Next, for $i=1,...,s$ the challenger computes a signature $\sigma_{i}=\textbf{Evaluate}(pk,\tau,f_{i},\overrightarrow{\sigma})$ of $f_{i}(\overrightarrow{\mathbf{m}}_{b})$. It sends the $(\tau,\sigma_{1},...,\sigma_{s})$ to $\mathcal{A}$. Note that the functions $f_{1},...,f_{s}$ can be output adaptively after $\overrightarrow{\mathbf{m}}_{0}$, $\overrightarrow{\mathbf{m}}_{1}$ are output.
Output: $\mathcal{A}$ output a bit $b^{\prime}$. The adversary $\mathcal{A}$ wins the game if $b^{\prime}=b$. The advantage of $\mathcal{A}$ is the probability that $\mathcal{A}$ wins the game.
We say that the linearly homomorphic signature scheme is $s$-weakly context hiding if the attacker cannot win the privacy game after seeing at most $s$ signatures derived from $\overrightarrow{\mathbf{m}}_{0}$, $\overrightarrow{\mathbf{m}}_{1}$.
2.3 Entropy and Statistical Distance
Definition 6.
(Statistical distance, [30]) Let $M\subset\mathbb{R}^{n}$ be a finite or countable set, and let $X$ and $Y$ be two discrete random variables taking values in $M$. The statistical distance between $X$ and $Y$ is defined as:
$\bigtriangleup(X,Y)=\frac{1}{2}\sum_{a\in M}|P\{X=a\}-P\{Y=a\}|.$
When the statistical distance between two distributions is less than a negligible value, we say that the two distributions are statistically close.
Definition 7.
([26], Min-entropy) For a random variable $X$, its min-entropy is defined as:
$H_{\infty}(X)=-\log\left(\max_{x\in X}\Pr[X=x]\right).$
The average conditional min-entropy of a random variable $X$ conditional on a correlated variable $Y$ is defined as:
$H_{\infty}(X|Y)=-\log\left(\mathbb{E}_{y\in Y}\left\{\max_{x\in X}\Pr[X=x|Y=y]\right\}\right).$
The optimal probability of an unbounded adversary guessing $X$ given the correlated value $Y$ is $2^{-H_{\infty}(X|Y)}$[26].
Lemma 1.
([26]) Let $X$, $Y$ be arbitrarily random variables where the support of $Y$ lies in $\mathcal{Y}$, Then
$H_{\infty}(X|Y)>H_{\infty}(X)-\log(|\mathcal{Y}|).$
2.4 Background on Lattices and Hard Problems
Definition 8.
(Lattice, [30] Let $\Lambda\subset\mathbb{R}^{n}$ be a non-empty subset. $\Lambda$ is called a lattice if:
(1) it is an additive subgroup of $\mathbb{R}^{n}$;
(2) there exists a positive constant $\lambda=\lambda(\Lambda)>0$ such that
$\mathrm{min}\{\parallel\mathbf{x}\parallel|\mathbf{x}\in\Lambda,\mathbf{x}\neq 0\}=\lambda.$
$\lambda$ is called the minimum distance.
A full-rank $n$-dimensional lattice can also be expressed as a linear combination of a set of basis vectors $B=\{\mathbf{b}_{1},...,\mathbf{b}_{n}\}\subset\mathbb{R}^{n}$:
$\Lambda=\mathcal{L}(B)=\{B\cdot\mathbf{x}=\sum_{i=1}^{n}x_{i}\mathbf{b}_{i}|x=(x_{1},...,x_{n})^{\top}\in\mathbb{Z}^{n}\}.$
We call $\Lambda^{*}$ the dual lattice of $\Lambda$ if
$\Lambda^{*}=\left\{\mathbf{y}\in\mathbb{R}^{n}\mid\langle\mathbf{y},\mathbf{x}\rangle\in\mathbb{Z},\quad\text{for all }\mathbf{x}\in\Lambda\right\}.$
Definition 9.
( $q$-ary lattices, [15]) Let $A\in\mathbb{Z}_{q}^{n\times m}$, $u\in\mathbb{Z}^{n}$. The following two $q$-ary lattices are defined as:
(1)$\Lambda_{q}^{\bot}=\{\mathbf{x}\in\mathbb{Z}^{m}|A\cdot\mathbf{x}\equiv 0(\mathrm{mod}q)\}$
(2)$\Lambda_{q}^{\mathbf{u}}=\{\mathbf{y}\in\mathbb{Z}^{m}|A\cdot\mathbf{y}\equiv\mathbf{u}(\mathrm{mod}q)\}$
The set $\Lambda_{q}^{\mathbf{u}}$ is a coset of $\Lambda_{q}^{\bot}$ since $\Lambda_{q}^{\mathbf{u}}=\Lambda_{q}^{\bot}+\mathbf{t}$ for any $\mathbf{t}$ such that $A\cdot\mathbf{t}\equiv\mathbf{u}(\mathrm{mod}q)$.
Definition 10.
(Short integer solution, [31]) Let $n$, $m$, $q$ be positive integers, with $m=\mathrm{poly(}n)$. Let $A\in\mathbb{Z}_{q}^{n\times m}$ be a uniformly distributed random matrix over $\mathbb{Z}_{q}$, and let $\beta\in\mathbb{R}$ such that $0<\beta<q$. The SIS problem is to find a short integer solution $\mathbf{x}$ satisfying the following condition:
$A\cdot\mathbf{x}\equiv 0(\mathrm{mod}q),\quad\mathrm{and}\quad\mathbf{x}\neq 0,\parallel\mathbf{x}\parallel\leq\beta.$
We denote the SIS problem with parameters $\textbf{SIS}_{q,n,m,\beta}$ or $\textbf{SIS}_{q,\beta}$.
Theorem 1.
(Worst-case to average-case reduction, [32]) In $\mathbb{Z}_{q}^{n\times m}$, $m=poly(n)$, $q\geq\beta\omega(\sqrt{n\log n})$, and $\beta=poly(n)$, if there is a polynomial time algorithm that solves $\textbf{SIS}_{q,\beta}$ with non-negligible probability, then there is a polynomial time algorithm that finds short non-zero vectors, which are only a $\lambda\geq\widetilde{\mathcal{O}}(\sqrt{n})$ factor longer than the shortest vector, in all lattices of dimension $n$.
2.5 Discrete Gaussian Distributions, Lattice Trapdoor Algorithms, and Sampling Algorithms
Definition 11.
(Discrete Gaussian Distribution, [30]) Let $s$ be a positive real number and $\mathbf{c}\in\mathbb{R}^{n}$ be a vector. The Gaussian function centered at $\mathbf{c}$ with parameter $s$ is defined as: $\rho_{s,\mathbf{c}}(\mathbf{x})=e^{-\frac{\pi}{s^{2}}\parallel\mathbf{x}-\mathbf{c}\parallel^{2}}.$ The discrete Gaussian measure $\mathcal{D}_{\Lambda,s,\mathbf{c}}$ defined on the lattice $\Lambda$ is given by:
$\mathcal{D}_{\Lambda,s,\mathbf{c}}(\mathbf{x})=\frac{\rho_{s,\mathbf{c}}(\mathbf{x})}{\rho_{s,\mathbf{c}}(\Lambda)},$
where $\rho_{s,\mathbf{c}}(\Lambda)=\sum_{\mathbf{x}\in\Lambda}\rho_{s,\mathbf{c}}(x)$.
Micciancio and Goldwasser [33] proved that a full-rank set $S$ in the lattice $\Lambda$ can be transformed into a basis $T$ such that both have similarly low Gram-Schmidt norms.
Lemma 2.
([33], Lemma 7.1) Let $\Lambda$ be an $n$-dimensional lattice. There is a deterministic polynomial-time algorithm that, given an arbitrary basis of and a full-rank set $S=\{s_{1},...,s_{n}\}$ in $\Lambda$, returns a basis $T$ of $\Lambda$ satisfying
$\parallel\widetilde{T}\parallel\leq\parallel\widetilde{S}\parallel\text{and}\parallel T\parallel\leq\parallel S\parallel\frac{\sqrt{n}}{2}.$
Theorem 2.
[…]
$\parallel\widetilde{T_{A}}\parallel\leq\mathcal{O}(\sqrt{h\log q})\quad\text{and}\quad\parallel T_{A}\parallel\leq\mathcal{O}(h\log q).$
Lemma 3.
(Sampling from Discrete Gaussian Distribution, [32]) Let $q\geq 2$ , $A\in\mathbb{Z}_{q}^{n\times m}$ and let $T_{A}$ be basis for $\Lambda_{q}^{\bot}(A)$ and $s\geq\widetilde{T_{A}}\cdot\omega(\sqrt{\log m})$. Then for $\mathbf{c}\in\mathbb{R}^{n}$ and $\mathbf{u}\in\mathbb{Z}^{n}$:
  1. 1.
    There is a probabilistic polynomial-time algorithm SampleGaussian$(A,T_{A},s,\mathbf{c})$ that outputs $\mathbf{x}\in\Lambda_{q}^{\bot}(A)$ drawn from a distribution statistically close to $\mathcal{D}_{\Lambda_{q}^{\bot}(A),s,\mathbf{c}}$
  2. 2.
    There is a probabilistic polynomial-time algorithm SamplePre$(A,T_{A},\mathbf{u},s)$ that outputs $\mathbf{x}\in\Lambda_{q}^{\mathbf{u}}(A)$ sampled from a distribution statistically close to $\mathcal{D}_{\Lambda_{q}^{\mathbf{u}}(A),s}$, whenever $\Lambda_{q}^{\mathbf{u}}(A)$ is not empty.
When $s\geq\omega(\sqrt{\log n})$, we denote the Gaussian sampling algorithm over the integer lattice $\mathbb{Z}^{n}$ as $\textbf{SampleDom}(1^{n},s)$. That is, when $\mathbf{x}\stackrel{{\scriptstyle\$}}{{\leftarrow}}\textbf{SampleDom}(1^{n},s)$, $\mathbf{x}$ is statistically close to the distribution $\mathcal{D}_{\mathbb{Z}^{n},s}$. And these preimages have a conditional min-entropy of $\omega(\log n)$[32].
Lemma 4.
([32]) Let $n$ and $h$ be positive integers, and $q$ be a prime, such that $n\geq 2h\lg q$. Then for all but a $2q^{-h}$ fraction of all $A\in\mathbb{Z}_{q}^{h\times n}$ and for any $s\geq\omega(\sqrt{\log n})$, the distribution of the $\alpha=A\cdot\mathbf{x}(\mathrm{mod}q)$ is statistically close to uniform over $\mathbb{Z}_{q}^{h}$, where $\mathbf{x}\stackrel{{\scriptstyle\$}}{{\leftarrow}}\textbf{SampleDom}(1^{n},s)$.
Lemma 5.
([34]) Let $\Lambda$ be an $n$-dimensional lattice, and $T$ be a basis of the lattice $\Lambda$. If $s\geq\parallel\widetilde{T}\parallel\cdot\omega(\sqrt{\log n})$, then for any $\mathbf{c}\in\mathbb{R}^{n}$, we have:
$\mathrm{Pr}\{\parallel\mathbf{x}-\mathbf{c}\parallel>s\sqrt{n}:\mathbf{x}{\leftarrow}\mathcal{D}_{\Lambda,s,\mathbf{c}}\}\leq\mathrm{negl}(n).$
Definition 12.
( Smoothing parameter[34]) For any $n$-dimensional lattice $\Lambda$ and any given $\epsilon>0$, the smoothing parameter of the lattice is defined as
$\eta_{\epsilon}(\Lambda)=\min\left\{s>0\mid\rho_{\frac{1}{s}}(\Lambda^{*})<1+\epsilon\right\}.$
Lemma 6.
([32]) Let $q\geq 3$, $h$ and $n$ be positive integers satisfying $n\geq 2h\lg q$. Then there exists a negligible function $\epsilon(n)$ such that $\eta_{\epsilon}(\Lambda_{q}^{\bot}(A))<\omega(\sqrt{\log n})$ for all but at most a $q^{-h}$ fraction of $A$ in the $\mathbb{Z}_{q}^{h\times n}$
Lemma 7.
([43]) For any $n$-dimensional lattice $\Lambda$, $\mathbf{c}\in\text{span}(\Lambda)$, real number $\epsilon\in(0,1)$, and $s\geq\eta_{\epsilon}(\Lambda)$, it holds that
$\underset{\mathbf{x}\sim D_{\Lambda,s,\mathbf{c}}}{\Pr}\left[\|\mathbf{x}-\mathbf{c}\|>s\sqrt{n}\right]\leq\frac{1+\epsilon}{1-\epsilon}\cdot 2^{-n}$
Agrawal et al. [24] noted that the algorithm ExtBasis used in the lattice basis delegation technique by Cash et al.[17] generates a new lattice with a dimension larger than the original one. Therefore, if this lattice basis delegation algorithm is used multiple times, the resulting lattice will have a very large dimension, significantly reducing the efficiency of lattice-based cryptographic schemes designed using this technique. To address this issue, Agrawal et al. developed a new lattice basis delegation technique that maintains a constant dimension during the delegation process. This technique requires the use of a so-called $\mathbb{Z}_{q}$-invertible matrix. As follows, we present the conclusion of Agrawal et al.[24].
Definition 13.
([24]) (1) We say that a matrix $R$ is $\mathbb{Z}_{q}$-invertible, if this matrix $R$ is invertible in $\mathbb{Z}_{q}^{n\times n}.$ (2) Let $s=\sqrt{h\log q}\cdot\omega(\sqrt{\log n})$. We let $\mathcal{D}_{n\times n}$ denote the distribution on matrices in $\mathbb{Z}^{n\times n}$ defined as $(\mathcal{D}_{(\mathbb{Z}^{n},s)})^{m}$ conditioned on the resulting matrix being $\mathbb{Z}_{q}$-invertible.
Theorem 3.
([24]) There exists a PPT algorithm BasisDel$(A,R,T_{A},s)$. Its inputs are: a matrix $A\in\mathbb{Z}_{q}^{h\times n}$ of rank $h$, a $\mathbb{Z}_{q}$-invertible matrix $R$ sampled from $\mathcal{D}_{n\times n}$, $T_{A}$ is a basis of $\Lambda_{q}^{\bot}(A)$, $s>\parallel\tilde{T}_{A}\parallel n\omega(\log^{2}n)$. The output of this algorithm is a basis $T_{B}$ of the lattice $\Lambda_{q}^{\bot}(B)$, where $B=AR^{-1}$, and $T_{B}<s/\omega(\sqrt{\log n})$.
Inspired by [24], we take a special case of the above algorithm BasisDel$(A,R,T_{A},s)$ that is, replace the input $R$ with an orthogonal matrix $H\in\mathbb{Z}^{n\times n}$. We can obtain a new algorithm NewBasis$(A,H,T_{A})$. Before presenting our algorithm, we first provide a necessary lemma.
Lemma 8.
Let $A\in\mathbb{R}^{n\times n}$ be an $n$-dimensional full-rank matrix, and $H\in\mathbb{R}^{n\times n}$ be an orthogonal matrix. Then,
$\parallel HA\parallel=\parallel A\parallel\quad\text{and}\quad\parallel\widetilde{HA}\parallel=\parallel\widetilde{A}\parallel.$
Proof.
(1) Suppose $A=(\mathbf{a}_{1},...,\mathbf{a}_{n})$, then $HA=(H\mathbf{a}_{1},...,H\mathbf{a}_{n})$. For any $i=1,...,n$ , we have
$\parallel H\mathbf{a}_{i}\parallel^{2}=(H\mathbf{a}_{i})^{T}(H\mathbf{a}_{i})=\mathbf{a}_{i}^{T}H^{T}H\mathbf{a}_{i}=\mathbf{a}_{i}^{T}\mathbf{a}_{i}=\parallel\mathbf{a}_{i}\parallel^{2}.$
Therefore, $\parallel HA\parallel=\parallel A\parallel.$
(2) Let $\widetilde{A}=(\widetilde{\mathbf{a}}_{1},...,\widetilde{\mathbf{a}}_{n}),$ where the vector group $\widetilde{\mathbf{a}}_{1},...,\widetilde{\mathbf{a}}_{n}$ is obtained by Gram-Schmidt orthogonalization of the vector group $\widetilde{\mathbf{a}}_{1},...,\widetilde{\mathbf{a}}_{n}$. Then we have $\widetilde{HA}=(H\widetilde{\mathbf{a}}_{1},...,H\widetilde{\mathbf{a}}_{n})$. The reasons are as follows.
We will prove this by mathematical induction. Given vectors $\mathbf{a}_{1},...,\mathbf{a}_{n}$, according to the definition of Gram-Schmidt orthogonalization:
$\widetilde{\mathbf{a}}_{1}=\mathbf{a}_{1},\quad\widetilde{\mathbf{a}}_{j}=\mathbf{a}_{j}-\sum_{i=1}^{j-1}\frac{\langle\mathbf{a}_{j},\widetilde{\mathbf{a}}_{i}\rangle}{\langle\widetilde{\mathbf{a}}_{i},\widetilde{\mathbf{a}}_{i}\rangle}\widetilde{\mathbf{a}}_{i}\quad j=2,...,n.$
Let $(\widetilde{HA})_{i}$ denote the $i$-th column vector of the matrix $\widetilde{HA}$. Since $\widetilde{\mathbf{a}}_{1}=\mathbf{a}_{1}$, we have $H\widetilde{\mathbf{a}}_{1}=H\mathbf{a}_{1}$. Moreover, the first column vector of $\widetilde{HA}$ is also $H\mathbf{a}_{1}$. Therefore, the first column vector of $\widetilde{HA}$ is identical to that of $H\widetilde{\mathbf{a}}_{1}$.
Assume that for $j=2,...,n-1$, we have $(\widetilde{HA})_{j}=H\widetilde{\mathbf{a}}_{j}$. We can easily prove that $(\widetilde{HA})_{n}=H\widetilde{\mathbf{a}}_{n},$ since
[…]
Therefore, $\widetilde{HA}=(H\widetilde{\mathbf{a}}_{1},...,H\widetilde{\mathbf{a}}_{n})$, and consequently, $\parallel\widetilde{HA}\parallel=\parallel\widetilde{A}\parallel.$
Theorem 4.
Let $q\geq 3$ be odd, $h$ be a positive integer, and let $n:=\lceil 6h\log q\rceil.$ There exists a deterministic polynomial-time algorithm denoted as NewBasis$(A,H,T_{A})$ Its inputs are: a matrix $A\in\mathbb{Z}_{q}^{h\times n}$ of rank $h$, a orthogonal matrix $H\in\mathbb{Z}^{n\times n}$, and $T_{A}$ which is a basis of $\Lambda_{q}^{\bot}(A)$. The output of this algorithm is a basis $T_{B}$ of the lattice $\Lambda_{q}^{\bot}(B)$, where $B=AH^{\top}$, and $\parallel\widetilde{T}_{B}\parallel<\parallel\widetilde{T}_{A}\parallel$, $\parallel T_{B}\parallel<\parallel T_{A}\parallel\sqrt{n}/2$
Proof.
  1. (1)
    According to Theorem 2, run the algorithm $\textbf{TrapGen}(q,h,n)$ to obtain matrix $A\in\mathbb{Z}_{q}^{h\times n}$ and $T_{A}$, where $T_{A}$ is the basis of $\Lambda_{q}^{\bot}(A)$. Take $A$ and $T_{A}$ as the inputs of the algorithm.
  2. (2)
    Select an orthogonal matrix $H$ from $\mathbb{Z}^{n\times n}$. For example, one can choose a diagonal matrix with diagonal entries equal to ±$\pm 1$. It is evident that such a matrix is an orthogonal matrix.
  3. (3)
    We regard $T=HT_{A}$ as a set of linearly independent vectors of the lattice $\Lambda_{q}^{\bot}(B)$, where $B=AH^{\top}$. According to Lemma 2 and Lemma 8, by inputting $T$ and an arbitrary basis of the lattice $\Lambda_{q}^{\bot}(B)$, a basis $T_{B}$ of the lattice $\Lambda_{q}^{\bot}(B)$ can be obtained in polynomial time, satisfying $\parallel\widetilde{T}_{B}\parallel<\parallel\widetilde{HT}_{A}\parallel=\parallel\widetilde{T_{A}}\parallel,\parallel T_{B}\parallel<\parallel HT_{A}\parallel\sqrt{n}/2=\parallel T_{A}\parallel\sqrt{n}/2$.
Remark 2: for any $q$-ary lattice, a basis can be found in polynomial time without any restrictions on the length of the basis vectors.
2.6 Statistical Properties Related to Discrete Gaussian Distributions
Theorem 5.
([14]) Let $\Lambda$ be an $n$-dimensional lattice, $A\in\mathbb{Z}^{s\times k}$, and $T\in\mathbb{Z}^{k\times n}$. $X$ is a random matrix of size $k\times n$ where the rows are independent, and the $i$-th row follows the distribution $\mathcal{D}_{\Lambda+\mathbf{t}_{i},\sigma}$, with $\mathbf{t}_{i}$ being the $i$-th row of the matrix $T$. If $\sigma>\mu(A)\eta_{\epsilon}(\Lambda)$ where $\epsilon$ is some negligible value and $\mu(A)=\sqrt{\mathrm{det}(AA^{\top})}$. Then the random variable $AX$ is statistically close to a distribution parameterized by $\Lambda$, $\sigma$, $A$, $AT$.
This theorem shows that for any matrix $A\in\mathbb{Z}^{s\times k}$ and sufficiently large $\sigma$, the random variable $AX$ is statistically close to a distribution parameterized by $\Lambda,\sigma,A$ ,$AT$. Furthermore, the dependence on the matrix $T$ is only reflected through the quantity $AT$. Therefore, if $X_{1}$ and $X_{2}$ are random variables defined in the same way as $X$, but using matrices $T_{1}$ and $T_{2}$, then when $AT_{1}=AT_{2}$, $AX_{1}$ and $AX_{2}$ are statistically close.
Lemma 9.
([14]) $\boldsymbol{t}_{i}\in\mathbb{Z}^{m}$ and $\boldsymbol{x}_{i}$ are mutually independent random variables sampled from a Gaussian distribution $D_{\boldsymbol{t}_{i}+\Lambda,\sigma}$ over $\boldsymbol{t}_{i}+\Lambda$ for $i=1,2,\cdots,k$ in which $\Lambda$ is a lattice and $\sigma\in\mathbb{R}$ is a parameter. Let $\boldsymbol{c}=(c_{1},\cdots,c_{k})\in\mathbb{Z}^{k}$ and $g=\mathrm{gcd}(c_{1},\cdots,c_{k})$, $\boldsymbol{t}=\sum_{i=1}^{k}c_{i}\boldsymbol{t}_{i}$. If $\sigma>\|\boldsymbol{c}\|\eta_{\epsilon}(\Lambda)$ for some negligible number $\epsilon$, then $\boldsymbol{z}=\sum_{i=1}^{k}c_{i}\boldsymbol{x}_{i}$ statistically closes to $D_{\boldsymbol{t}+g\Lambda,\|\boldsymbol{c}\|\sigma}$.
Lemma 10.
Let $H$ be an orthogonal matrix over $\mathbb{Z}^{n\times n}$, that is, $HH^{\top}=I_{n}$. If $\mathbf{x}\sim\mathcal{D}_{\mathbb{Z}^{n},\sigma}$, then $H\mathbf{x}\sim\mathcal{D}_{\mathbb{Z}^{n},\sigma}$.
Proof.
(1) Clearly, $H\cdot\mathbb{Z}^{n}=\{H\mathbf{x}|\mathbf{x}\in\mathbb{Z}^{n}\}$ is a lattice. Next, we prove that $\mathbb{Z}^{n}=H\cdot\mathbb{Z}^{n}$.
For any $\mathbf{x}\in\mathbb{Z}^{n}$, we have $\mathbf{x}=H\cdot(H^{\top}\cdot\mathbf{x})$. Since $H^{\top}\cdot\mathbf{x}\in\mathbb{Z}^{n}$, then $\mathbf{x}\in H\cdot\mathbb{Z}^{n}$. Conversely, for any $\mathbf{x}\in H\cdot\mathbb{Z}^{n}$, obviously $\mathbf{x}\in\mathbb{Z}^{n}$. Therefore, $\mathbb{Z}^{n}=H\cdot\mathbb{Z}^{n}$.
(2) Next, we prove that the probability densities of $\mathbf{x}$ and $H\mathbf{x}$ are the same. By the definition of the discrete Gaussian measure:
$\mathcal{D}_{\mathbb{Z}^{n},s}(\mathbf{x})=\frac{\rho_{s}(\mathbf{x})}{\rho_{s}(\mathbb{Z}^{n})}=\frac{e^{-\frac{\pi}{s^{2}}\|\mathbf{x}\|^{2}}}{\rho_{s}(\mathbb{Z}^{n})}=\frac{e^{-\frac{\pi}{s^{2}}\|H\mathbf{x}\|^{2}}}{\rho_{s}(H\cdot\mathbb{Z}^{n})}=\mathcal{D}_{H\cdot\mathbb{Z}^{n},s}(H\mathbf{x}).$
Therefore, $H\mathbf{x}\sim\mathcal{D}_{\mathbb{Z}^{n},s}$. ∎
Lemma 11.
Let $A\in\mathbb{Z}_{q}^{n\times m},s>0,\mathbf{u}\in\mathbb{Z}^{m}$, and $\Lambda=\Lambda_{q}^{\boldsymbol{u}}(A)$. If $\mathbf{x}$ is sampled from $\mathcal{D}_{\mathbb{Z}^{m},s}$ conditioned on $A\mathbf{x}\equiv\boldsymbol{u}\pmod{q}$, then the conditional distribution of $\mathbf{x}$ is $\mathcal{D}_{\Lambda,s}$.
Proof.
According to Definition 9, the discrete Gaussian distribution defined on $\Lambda$ is:
$\mathcal{D}_{\Lambda,s}(x)=\frac{\rho_{s}(\mathbf{x})}{\rho_{s}(\Lambda)},\quad\mathbf{x}\in\Lambda,\text{ where }\rho_{s}(\mathbf{x})=e^{-\frac{\pi}{s^{2}}\|\mathbf{x}\|^{2}}.$
Then, for $\mathbf{x}\sim\mathcal{D}_{\mathbb{Z}^{m},s}$ and $\mathbf{x}\in\Lambda$, it can be regarded as the following conditional distribution:
$\mathcal{D}_{\mathbb{Z}^{m},s}(\mathbf{x}\mid\mathbf{x}\in\Lambda)=\frac{\mathcal{D}_{\mathbb{Z}^{m},s}(\mathbf{x})}{\mathcal{D}_{\mathbb{Z}^{m},s}(\Lambda)}=\frac{\rho_{s}(x)/\rho_{s}(\mathbb{Z}^{m})}{\rho_{s}(\Lambda)/\rho_{s}(\mathbb{Z}^{m})}=\frac{\rho_{s}(\mathbf{x})}{\rho_{s}(\Lambda)}=\mathcal{D}_{\Lambda,s}(\mathbf{x}).$
Therefore, $\mathbf{x}\sim\mathcal{D}_{\Lambda,s}$.
Theorem 6.
Assume that $s_{1}\geq s_{2}\geq\eta_{\epsilon}(\mathbb{Z}^{n})$, where $\epsilon=\frac{1}{\mathrm{poly}(n)}$ is a sufficiently small value. If $|\frac{1}{s_{1}^{2}}-\frac{1}{s_{2}^{2}}|\leq\frac{\epsilon}{2\pi ns_{1}^{2}}$ and $(\frac{s_{2}}{s_{1}})^{n}\geq 1-\frac{\epsilon}{1-\epsilon}$, then the statistical distance between $D_{\mathbb{Z}^{n},s_{1}}$ and $D_{\mathbb{Z}^{n},s_{2}}$ is at most $3\epsilon$.
Proof.
By the Poisson summation formula, we have:
$\rho_{s}(\mathbb{Z}^{n})=\sum_{\mathbf{x}\in\mathbb{Z}^{n}}\exp\left(-\pi\frac{\|\mathbf{x}\|^{2}}{s^{2}}\right)=s^{n}\sum_{\mathbf{x}\in\mathbb{Z}^{n}}\exp\left(-\pi\|\mathbf{x}\|^{2}s^{2}\right).$
Since $s>\eta_{\epsilon}(\mathbb{Z}^{n})$ and the dual lattice of $\mathbb{Z}^{n}$ is itself, it follows that:
$\rho_{s}(\mathbb{Z}^{n})=s^{n}\sum_{\mathbf{x}\in\mathbb{Z}^{n}}\exp\left(-\pi\|\mathbf{x}\|^{2}s^{2}\right)=s^{n}\rho_{1/s}(\mathbb{Z}^{n})\leq s^{n}(1+\epsilon).$
Thus, we obtain the inequality:
$s_{i}^{n}\leq\rho_{s_{i}}(\mathbb{Z}^{n})\leq s_{i}^{n}(1+\epsilon).$
Consider the ratio $\frac{\rho_{s_{2}}(\mathbb{Z}^{n})}{\rho_{s_{1}}(\mathbb{Z}^{n})}$. Noting that $\frac{s_{2}}{s_{1}}\leq 1$, we have:
$\left(\frac{s_{2}}{s_{1}}\right)^{n}\frac{1}{1+\epsilon}\leq\frac{\rho_{s_{2}}(\mathbb{Z}^{n})}{\rho_{s_{1}}(\mathbb{Z}^{n})}\leq\frac{(1+\epsilon)s_{2}^{n}}{s_{1}^{n}}\leq 1+\epsilon.$
Let $U=\{\mathbf{x}\mid\|\mathbf{x}\|\leq 1\}$ and $B=\{\mathbf{x}\mid\|\mathbf{x}\|\leq s_{1}\sqrt{n}\}$. For sufficiently small $\epsilon$ (e.g., $\epsilon=1/\text{poly}(n)$), by Theorem 6 (for large $n$), we get:
$\Pr_{\mathbf{x}\sim\mathcal{D}_{\mathbb{Z}^{n},s_{1}}}\left[\|\mathbf{x}\|>s_{1}\sqrt{n}\right]=\frac{\rho_{s_{1}}(\mathbb{Z}^{n}\setminus s_{1}\sqrt{n}U)}{\rho_{s_{1}}(\mathbb{Z}^{n})}=\sum_{\mathbf{x}\notin B}D_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})\leq\frac{1+\epsilon}{1-\epsilon}\cdot 2^{-n}\leq\epsilon.$
Similarly,
$\sum_{\mathbf{x}\notin B}D_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})=\frac{\rho_{s_{2}}(\mathbb{Z}^{n}\setminus s_{1}\sqrt{n}U)}{\rho_{s_{2}}(\mathbb{Z}^{n})}\leq\frac{\rho_{s_{2}}(\mathbb{Z}^{n}\setminus s_{2}\sqrt{n}U)}{\rho_{s_{2}}(\mathbb{Z}^{n})}\leq\frac{1+\epsilon}{1-\epsilon}\cdot 2^{-n}\leq\epsilon.$
Thus,
$\sum_{\mathbf{x}\notin B}\left|\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})-\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})\right|\leq\sum_{\mathbf{x}\notin B}\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})+\sum_{\mathbf{x}\notin B}\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})\leq 2\epsilon.$
For $\mathbf{x}\in B$, consider the ratio:
$\frac{\rho_{s_{1}}(\mathbf{x})}{\rho_{s_{2}}(\mathbf{x})}=\exp\left(-\pi\|\mathbf{x}\|^{2}\left(\frac{1}{s_{1}^{2}}-\frac{1}{s_{2}^{2}}\right)\right).$
Notice that:
$\left|\pi\|\mathbf{x}\|^{2}\left(\frac{1}{s_{1}^{2}}-\frac{1}{s_{2}^{2}}\right)\right|\leq\pi s_{1}^{2}n\cdot\frac{\epsilon}{2\pi ns_{1}^{2}}=\frac{\epsilon}{2}.$
Since $|e^{z}-1|\leq 2|z|$ for $|z|\leq\frac{1}{2}$, we have:
$1-\epsilon\leq\frac{\rho_{s_{1}}(\mathbf{x})}{\rho_{s_{2}}(\mathbf{x})}\leq 1+\epsilon.$
For $\mathbf{x}\in B$, it follows that:
$\frac{\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})}{\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})}=\frac{\rho_{s_{1}}(\mathbf{x})}{\rho_{s_{2}}(\mathbf{x})}\cdot\frac{\rho_{s_{2}}(\mathbb{Z}^{n})}{\rho_{s_{1}}(\mathbb{Z}^{n})}\subset\left[\left(\frac{s_{2}}{s_{1}}\right)^{n}\frac{1-\epsilon}{1+\epsilon},(1+\epsilon)^{2}\right]\subset[1-3\epsilon,1+4\epsilon].$
Hence,
$\left|\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})-\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})\right|\leq 4\epsilon\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x}),$
and therefore:
$\sum_{\mathbf{x}\in B}\left|\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})-\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})\right|\leq 4\epsilon\sum_{\mathbf{x}\in B}\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})\leq 4\epsilon.$
Combining these results, we calculate:
$\frac{1}{2}\sum_{\mathbf{x}\in\mathbb{Z}^{n}}$ $\left|\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})-\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})\right|$
$=\frac{1}{2}\sum_{\mathbf{x}\in B}\left|\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})-\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})\right|$
$\quad+\frac{1}{2}\sum_{\mathbf{x}\notin B}\left|\mathcal{D}_{\mathbb{Z}^{n},s_{1}}(\mathbf{x})-\mathcal{D}_{\mathbb{Z}^{n},s_{2}}(\mathbf{x})\right|$
$\leq\frac{1}{2}(4\epsilon+2\epsilon)$
$\leq 3\epsilon.$
Corollary 7.
Let $k=\text{poly}(n)$ be even, $q=\text{poly}(n)\geq(nk)^{3}$, and $V=\sqrt{2n\log q}\log n\geq\omega(\log n)$. Then $\mathcal{D}_{\mathbb{Z}^{n},V}$ and $\mathcal{D}_{\mathbb{Z}^{n},\sqrt{\frac{k\pm 2}{k}}V}$ are statistically indistinguishable.
Proof.
Let $s_{1}=V$ and $s_{2}=\sqrt{\frac{k-2}{k}}V$. We proceed to show that $\mathcal{D}_{\mathbb{Z}^{n},s_{1}}$ and $\mathcal{D}_{\mathbb{Z}^{n},s_{2}}$ are statistically indistinguishable.
By Lemma 6 and the construction of $V$, there exists a sufficiently small $\epsilon$ such that $s_{1}>\eta_{\epsilon}(\mathbb{Z}^{n})$ and $s_{2}>\eta_{\epsilon}(\mathbb{Z}^{n})$. For convenience, let $\epsilon=\frac{1}{n^{d}}$ where $d>0$ and $k=n^{d+2}$. When $n$ is sufficiently large, we have:
$\left|\frac{1}{s_{1}^{2}}-\frac{1}{s_{2}^{2}}\right|=\left|\frac{1}{V^{2}}-\frac{1}{\frac{k-2}{k}V^{2}}\right|=\frac{2}{(k-2)V^{2}}\leq\frac{1}{2\pi n^{d+1}V^{2}}=\frac{\epsilon}{2\pi ns_{1}^{2}}.$
Next, applying Bernoulli’s inequality: for any $x\in(0,1)$ and $y\geq 1$, $(1-x)^{y}\geq 1-xy$. Thus,
$\left(\frac{s_{2}}{s_{1}}\right)^{n}$ $=\left(1-\frac{2}{k}\right)^{n/2}=\left(1-\frac{2}{n^{d+2}}\right)^{n/2}$ (1)
$\geq 1-\frac{2}{n^{d+2}}\cdot\frac{n}{2}=1-\frac{1}{n^{d+1}}$
$\geq 1-\frac{1}{n^{d}-1}=1-\frac{1}{\frac{1}{\epsilon}-1}=1-\frac{\epsilon}{1-\epsilon}.$
By Theorem 6, $\mathcal{D}_{\mathbb{Z}^{n},V}$ and $\mathcal{D}_{\mathbb{Z}^{n},\sqrt{\frac{k-2}{k}}V}$ are statistically indistinguishable.
Following identical steps, we can show that $\mathcal{D}_{\mathbb{Z}^{n},V}$ and $\mathcal{D}_{\mathbb{Z}^{n},\sqrt{\frac{k+2}{k}}V}$ are also statistically indistinguishable. This completes the proof. ∎
3 New linearly homomorphic signature Scheme
Before introducing the specific construction methods, it is first necessary to propose a decomposition algorithm named Decompose. The detailed content is as follows:
Theorem 8.
Let $k$ be an odd integer with $k=\mathrm{poly}(n)$. Define the sets
$\mathcal{X}$ =$=\{x\in\mathbb{F}_{2}^{2k}\mid\|x\|=k-1\}\sqcup\{0\},$
$\mathcal{Y}$ =$=\{y\in\mathbb{F}_{2}^{2k}\mid\|y\|=k\}\sqcup\{0\},$
$\mathcal{Z}$ =$=\{z\in\mathbb{F}_{2}^{2k}\mid\|z\|=k+1\}\sqcup\{0\}.$
There exists a polynomial-time algorithm $\textbf{Decompose}(\mathbf{w})$ that, given an input $\mathbf{w}\in\mathbb{F}_{2}^{2k}$, outputs a pair of vectors $(\mathbf{u},\mathbf{v})$ with $\mathbf{u}\in\mathcal{X}$ and $\mathbf{v}\in\mathcal{Y}$ or $\mathcal{Z}$.
Proof.
For any $\mathbf{w}\in\mathbb{F}_{2}^{2k}$, let $\|\mathbf{w}\|=t$. The algorithm proceeds as follows:
Case 1: $t$ is odd
(1) If $t=k$, set $\mathbf{u}=0$ and $\mathbf{v}=\mathbf{w}$. Clearly, $\mathbf{w}=\mathbf{u}+\mathbf{v}$ with $\mathbf{u}\in\mathcal{X}$ and $\mathbf{v}\in\mathcal{Y}$, so the algorithm outputs $(\mathbf{u},\mathbf{v})$.
(2) If $1\leq t<k$, suppose the coordinates of $\mathbf{w}$ equal to 1 are $i_{1},i_{2},\dots,i_{t}$ where $1\leq i_{s}\leq 2k$ for $s=1,\dots,t$.
  • $\bullet$
    Compute a pair $(n,m)$ such that $t=n+m$ and $n-m=1$;
  • $\bullet$
    Construct $\mathbf{v}$: Select $n$ coordinates from $i_{1},i_{2},\dots,i_{t}$ and set them to 1. Choose $k-n$ coordinates from the remaining $2k-t$ coordinates and set them to 1. All other coordinates are 0. Denote this vector as $\mathbf{v}$;
  • $\bullet$
    Construct $\mathbf{u}$: Set the remaining $m$ coordinates in $i_{1},i_{2},\dots,i_{t}$ (not selected for $\mathbf{v}$) to 1, and set the $n$ selected coordinates to 0. All other coordinates match $\mathbf{v}$. Denote this vector as $\mathbf{u}$.
Verification: Clearly, $\mathbf{w}=\mathbf{u}+\mathbf{v}$. Moreover, $\|\mathbf{u}\|=m+k-n=k-(n-m)=k-1$ and $\|\mathbf{v}\|=n+(k-n)=k$, so $\mathbf{u}\in\mathcal{X}$ and $\mathbf{v}\in\mathcal{Y}$. The algorithm outputs $(\mathbf{u},\mathbf{v})$.
(3) If $k<t\leq 2k-1$, suppose the coordinates of $\mathbf{w}$ equal to 1 are $i_{1},i_{2},\dots,i_{t}$.
  • $\bullet$
    Compute a pair $(n,m)$ such that $t=n+m$ and $n-m=1$;
  • $\bullet$
    Construct $\mathbf{v}$: Select $n$ coordinates from $i_{1},i_{2},\dots,i_{t}$ and set them to 1. Choose $\frac{2k-1-t}{2}$ coordinates from the remaining $2k-t$ coordinates and set them to 1. All other coordinates are 0. Denote this vector as $\mathbf{v}$;
  • $\bullet$
    Construct $\mathbf{u}$: Set the remaining $m$ coordinates in $i_{1},i_{2},\dots,i_{t}$ (not selected for $\mathbf{v}$) to 1, and set the $n$ selected coordinates to 0. All other coordinates match $\mathbf{v}$. Denote this vector as $\mathbf{u}$.
Verification: Clearly, $\mathbf{w}=\mathbf{u}+\mathbf{v}$. Moreover,
$\|\mathbf{u}\|$ $=m+\frac{2k-1-t}{2}=m+\frac{2k-1-(n+m)}{2}=k-1,$
$\|\mathbf{v}\|$ $=n+\frac{2k-1-t}{2}=n+\frac{2k-1-(n+m)}{2}=k,$
so $\mathbf{u}\in\mathcal{X}$ and $\mathbf{v}\in\mathcal{Y}$. The algorithm outputs $(\mathbf{u},\mathbf{v})$.
Case 2: $t$ is even
(1) If $t=0$, set $\mathbf{u}=\mathbf{w}$ and $\mathbf{v}=\mathbf{w}$. This trivially satisfies the output conditions, so the algorithm outputs $(\mathbf{u},\mathbf{v})$.
(2) If $t=2k$, randomly select a non-zero vector $\mathbf{u}\in\mathcal{X}$ and set $\mathbf{v}=\mathbf{w}+\mathbf{u}$. Then $\mathbf{w}=\mathbf{u}+\mathbf{v}$ and $\|\mathbf{v}\|=2k-(k-1)=k+1$, so $\mathbf{v}\in\mathcal{Z}$. The algorithm outputs $(\mathbf{u},\mathbf{v})$.
(3) If $t=k-1$, set $\mathbf{u}=\mathbf{w}$ and $\mathbf{v}=0$. This satisfies the output conditions, so the algorithm outputs $(\mathbf{u},\mathbf{v})$.
(4) If $t=k+1$, set $\mathbf{u}=0$ and $\mathbf{v}=\mathbf{w}$. This satisfies the output conditions, so the algorithm outputs $(\mathbf{u},\mathbf{v})$.
(5) If $12\leq t<k-$ 1, suppose the coordinates of $\mathbf{w}$ equal to 1 are $i_{1},i_{2},\dots,i_{t}$.
  • $\bullet$
    Compute a pair $(n,m)$ such that $t=n+m$ and $n-m=2$;
  • $\bullet$
    Construct $\mathbf{v}$: Select $n$ coordinates from $i_{1},i_{2},\dots,i_{t}$ and set them to 1. Choose $k+1-n$ coordinates from the remaining $2k-t$ coordinates and set them to 1. All other coordinates are 0. Denote this vector as $\mathbf{v}$;
  • $\bullet$
    Construct $\mathbf{u}$: Set the remaining $m$ coordinates in $i_{1},i_{2},\dots,i_{t}$ (not selected for $\mathbf{v}$) to 1, and set the $n$ selected coordinates to 0. All other coordinates match $\mathbf{v}$. Denote this vector as $\mathbf{u}$.
Verification: Clearly, $\mathbf{w}=\mathbf{u}+\mathbf{v}$. Moreover, $\|\mathbf{u}\|=m+k+1-n=k+1-(n-m)=k-1$ and $\|\mathbf{v}\|=n+k+1-n=k+1$, so $\mathbf{u}\in\mathcal{X}$ and $\mathbf{v}\in\mathcal{Z}$. The algorithm outputs $(\mathbf{u},\mathbf{v})$.
(6) If $k+1<t\leq 2k-2$, suppose the coordinates of $\mathbf{w}$ equal to 1 are $i_{1},i_{2},\dots,i_{t}$.
  • $\bullet$
    Compute a pair $(n,m)$ such that $t=n+m$ and $n-m=2$;
  • $\bullet$
    Construct $\mathbf{v}$: Select $n$ coordinates from $i_{1},i_{2},\dots,i_{t}$ and set them to 1. Choose $\frac{2k-t}{2}$ coordinates from the remaining $2k-t$ coordinates and set them to 1. All other coordinates are 0. Denote this vector as $\mathbf{v}$;
  • $\bullet$
    Construct $\mathbf{u}$: Set the remaining $m$ coordinates in $i_{1},i_{2},\dots,i_{t}$ (not selected for $\mathbf{v}$) to 1, and set the $n$ selected coordinates to 0. All other coordinates match $\mathbf{v}$. Denote this vector as $\mathbf{u}$.
Verification: Clearly, $\mathbf{w}=\mathbf{u}+\mathbf{v}$. Moreover,
$\|\mathbf{u}\|$ $=m+\frac{2k-t}{2}=m+\frac{2k-(n+m)}{2}=k-1,$
$\|\mathbf{v}\|$ $=n+\frac{2k-t}{2}=n+\frac{2k-(n+m)}{2}=k+1,$
so $\mathbf{u}\in\mathcal{X}$ and $\mathbf{v}\in\mathcal{Z}$. The algorithm outputs $(\mathbf{u},\mathbf{v})$.
In all cases, the algorithm finds a valid decomposition $(\mathbf{u},\mathbf{v})$ in polynomial time. ∎
3.1 Construction
The algorithms TrapGen, NewBasis, SamplePre and Decompose are derived from Theorem 2, Theorem 4, Lemma 3, and Theorem 8 respectively. In this paper, the message space is $\mathcal{M}=\mathbb{F}_{2}^{k}$, and the signature space is $\Sigma=\mathbb{Z}^{n\times n}$. The definitions of sets $\mathcal{X}$, $\mathcal{Y}$, and $\mathcal{Z}$ are consistent with those in Theorem Theorem 8. Let $\langle f\rangle=(c_{1},...,c_{k})$ denote the linear function $f(\mathbf{m}_{1},...,\mathbf{m}_{k})=c_{1}\mathbf{m}_{1}+...+c_{k}\mathbf{m}_{k}$, where $\mathbf{m}_{i}\in\mathbb{F}_{2}^{k}$ and $c_{i}\in\{0,1\}$ for all $i$.
[…]
-Setup$(1^{n},k)$: Input $(n,k)$, where $n$ is the security parameter and $k=\mathsf{poly}(n)$ is the maximum size of the data set (the maximum number of linear operations), and $k$ is even. The parameters are generated as follows:
Let prime $q=\mathrm{poly}(n)$ such that $q\geq(nk)^{2}.$ Let $h=\lfloor\frac{n}{6\log q}\rfloor$, $V=\sqrt{2nk\log q}\cdot\log n$. Output public parameters $pp=(q,h,V)$.
-KeyGen($pp$): Input the public parameters $pp$. The algorithm generates a pair of public key and secret key as follows:
  1. (1)
    Compute $(A,T_{A})\longleftarrow\textbf{TrapGen}(q,h,n)$, where $A\in\mathbb{Z}_{q}^{h\times n}$ has rank $h$ and $T_{A}$ is a basis for the lattice $\Lambda_{q}^{\bot}(A)$.
  2. (2)
    Sample $\alpha_{1},\ldots,\alpha_{k}\stackrel{{\scriptstyle\$}}{{\leftarrow}}\mathbb{Z}_{q}^{h}$. We require that $\alpha_{1},\ldots,\alpha_{k}$ are linearly independent. If not, resampling should be carried out, and this process can be easily accomplished within polynomial time.
  3. (3)
    Output the public key $pk=(A,\alpha_{1},\ldots,\alpha_{k})$ and the secret key $sk=T_{A}$.
-Sign$(sk,\tau,\mathbf{m}_{i},i)$: Input the secret key $sk=T_{A}$, the tag $\tau\in\{0,1\}^{n}$, the data message $\mathbf{m}_{i}=(m_{i1},...,m_{ik})\in\mathbb{F}_{2}^{k}$ and the index $i$ indicating that $\mathbf{m}_{i}$ is the $i$-th data message in the data set. the signing algorithm proceeds as follows:
  1. (1)
    Let $H_{\tau}=\mathrm{diag}\{2\tau_{1}-1,2\tau_{2}-1,...,2\tau_{n}-1\}$, then $H_{\tau}$ is an orthogonal matrix.
  2. (2)
    Compute $T_{B^{\tau}}{\leftarrow}\textbf{NewBasis}(A,H_{\tau},T_{A})$, where $T_{B^{\tau}}$ is a basis for $\Lambda_{q}^{\bot}(B^{\tau})$ and $B^{\tau}=AH_{\tau}^{\top}$. Let $\Lambda^{\tau}=\Lambda_{q}^{\bot}(B^{\tau})$.
  3. (3)
    Compute $(\mathbf{u}_{i},\mathbf{v}_{i}){\leftarrow}\textbf{Decompose}(\mathbf{m}_{i})$, where $\mathbf{m_{i}}=(m_{i1},...,m_{ik})=\mathbf{u}_{i}+\mathbf{v}_{i}$, $\mathbf{u}_{i}=(u_{i1},...,u_{ik})\in\mathcal{X}$, $\mathbf{v}_{i}=(v_{i1},...,v_{ik})\in\mathcal{Y}$ or $\mathcal{Z}$.
  4. (4)
    Check if either vector in $(\mathbf{u}_{i},\mathbf{v}_{i})$ is zero. If so, compute $\mathbf{t}_{i}=\sum_{j=1}^{k}m_{ij}\alpha_{j}$, and $\sigma_{i}\longleftarrow\textbf{SamplePre}(\Lambda^{\tau},T_{B^{\tau}},\mathbf{t}_{i},V)$. Otherwise, proceed to the next step.
  5. (5)
    Let $\mathbf{t}(\mathbf{u}_{i})=\sum_{j=1}^{k}u_{ij}\alpha_{j}$ , $\mathbf{t}(\mathbf{v}_{i})=\sum_{j=1}^{k}v_{ij}\alpha_{j}$. Compute
    $\sigma(\mathbf{u}_{i})\longleftarrow\textbf{SamplePre}(\Lambda^{\tau},T_{B^{\tau}},\mathbf{t}\mathbf{(}\mathbf{u}_{i}),V),$
    $\sigma(\mathbf{v}_{i})\longleftarrow\textbf{SamplePre}(\Lambda^{\tau},T_{B^{\tau}},\mathbf{t}(\mathbf{v}_{i}),V),$
    Let $\sigma_{i}=\sigma(\mathbf{u}_{i})+\sigma(\mathbf{v}_{i})$.
  6. (6)
    Output $\sigma_{i}$ as signature.
-Verify$(pk,\tau,\mathbf{m},\sigma,f)$: Input the public key $pk$, a binary vector $\tau\in\{0,1\}^{n}$, a message $\mathbf{m}\in\mathbb{Z}_{2}^{k}$, a signature $\sigma\in\mathbb{Z}^{n}$, and a linear function $f=(c_{1},\ldots,c_{k})$, where each $c_{i}\in\{0,1\}$. The message $\mathbf{m}$ is defined as $\mathbf{m}=f(\mathbf{m}_{1},\ldots,\mathbf{m}_{k})$. The verification process is as follows:
  1. (1)
    Based on the input $pk=(A,\alpha_{1},\ldots,\alpha_{k})$ and $\tau$, compute $H_{\tau}=\mathrm{diag}\{2\tau_{1}-1,2\tau_{2}-1,...,2\tau_{n}-1\}$ and $B^{\tau}=AH_{\tau}^{\top}$. compute $\mathbf{t}=\sum_{i=1}c_{i}\mathbf{t}_{i}$, where $\mathbf{t}_{i}=\sum_{j=1}^{k}m_{ij}\alpha_{j}$, $i=1,...,k$.
  2. (2)
    Output 1 if all of the following three conditions are satisfied; otherwise, output 0.
    1. 1)
      $\parallel\sigma\parallel\leq k\cdot V\cdot\sqrt{2kn}$;
    2. 2)
      $B^{\tau}\cdot\sigma(\mathrm{mod}q)=\mathbf{t}$.
-Evaluate$(pk,\tau,f,\overrightarrow{\sigma})$: Input the public key $pk$, $\tau\in\{0,1\}^{n}$, a set of signatures $\overrightarrow{\sigma}=(\sigma_{1},...,\sigma_{k})$, where $\sigma_{i}=\textbf{Sign}(sk,\tau,\mathbf{m},i)$ and the linear function $\langle f\rangle=(c_{1},...,c_{k})$. The evaluation algorithm outputs $\sigma=\sum_{i=1}^{k}c_{i}\sigma_{i}$ as the signature of $f(\mathbf{m}_{1},...,\mathbf{m}_{k})$.
Remark 3: In the vector decomposition of message $\mathbf{m}_{i}$, when one of the components is a zero vector, the proof approach for the related conclusions follows the same methodology as the case where both components are non-zero vectors. Therefore, when demonstrating the correctness, unforgeability, and privacy of the scheme, it suffices to consider the scenario where $\mathbf{m}_{i}$ is decomposed into two non-zero vectors.
The following proves that the above scheme satisfies correctness.
Theorem 9.
The above linear homomorphic signature scheme satisfies correctness with overwhelming probability.
Proof.
(1) Given $\sigma_{i}=\operatorname{Sign}(sk,\tau,\mathbf{m}_{i},i)$, let $f=\pi_{i}$ be the projection function, which can be viewed as a special type of linear function, that is, $\pi_{i}(\mathbf{m}_{1},\dots,\mathbf{m}_{k})=\mathbf{m}_{i}$. In this case, we check the following two verification conditions:
From Theorem 2 and Theorem 4, it follows that $\|\widetilde{T}_{B^{\tau}}\|\leq\|\widetilde{T}_{A}\|\leq\mathcal{O}(\sqrt{h\log q})$ with overwhelming probability. Since $V=k\sqrt{2n\log q}\cdot\log n$, it follows that
$\frac{V}{\|\widetilde{T}_{B^{\tau}}\|}\geq k\sqrt{\frac{2n}{h}}\cdot\log n\geq\sqrt{\log n},$
which means $V\geq\|\widetilde{T}_{B^{\tau}}\|\cdot\omega(\sqrt{\log n})$.
Therefore, by Lemma 5, we have $\|\sigma(\mathbf{u}_{i})\|\leq V\cdot\sqrt{n}$ and $\|\sigma(\mathbf{v}_{i})\|\leq V\cdot\sqrt{n}$ with overwhelming probability. Therefore,
$\|\sigma_{i}\|\leq 2V\cdot\sqrt{n}<k\cdot V\cdot\sqrt{2kn}.$
Since
$B^{\tau}\cdot\sigma_{i}\pmod{q}$ $=B^{\tau}\cdot(\sigma(\mathbf{u}_{i})+\sigma(\mathbf{v}_{i}))\pmod{q}$
$=\mathbf{t}(\mathbf{u}_{i})+\mathbf{t}(\mathbf{v}_{i})$
$=\sum_{j=1}u_{ij}\alpha_{j}+\sum_{j=1}v_{ij}\alpha_{j}$
$=\sum_{j=1}(u_{ij}+v_{ij})\alpha_{j}$
$=\sum_{j=1}m_{ij}\alpha_{j}=\mathbf{t}_{i},$
thus, in the case where $f=\pi_{i}$, the scheme satisfies correctness.
(2) Given $(\mathbf{m}_{1},...,\mathbf{m}_{k})\in(\mathcal{M})^{k}$, $\overrightarrow{\sigma}=(\sigma_{1},...,\sigma_{k})$, where $\sigma_{i}=\textbf{Sign}(sk,\tau,\mathbf{m}_{i},i)$. Let $f=(c_{1},...,c_{k})$, where $c_{i}\in\{0,1\}$. According to the definition of the algorithm $\textbf{Evaluate}(pk,\tau,f,\vec{\sigma})$, $\sigma=\sum_{i=1}^{k}c_{i}\sigma_{i}$ is the signature of $f(\mathbf{m}_{1},...,\mathbf{m}_{k})$. In this case, we will verify that $\sigma$ satisfies the following two verification conditions:
Since $\parallel\sigma_{i}\parallel\leq 2V\cdot\sqrt{n}$ holds with overwhelming probability, and $|c_{i}|\leq 1$, we have
$\parallel\sigma\parallel=\parallel c_{1}\sigma_{1}+...+c_{k}\sigma_{k}\parallel\leq k\cdot\mathrm{max}_{1\leq i\leq k}\parallel\sigma_{i}\parallel\leq 2k\cdot V\cdot\sqrt{n}\leq k\cdot V\cdot\sqrt{2kn}$
Since $\sigma_{i}=\text{Sign}(sk,\tau,\mathbf{m}_{i},i)$, we have $B^{\tau}\cdot\sigma_{i}\pmod{q}=\mathbf{t}_{i}$. Therefore, $B^{\tau}\cdot c_{i}\sigma_{i}\pmod{q}=c_{i}\mathbf{t}_{i}$. By the linearity of modular arithmetic, we can derive:
$B^{\tau}\cdot\sigma\pmod{q}$ $=B^{\tau}\cdot\left(\sum_{i=1}^{k}c_{i}\sigma_{i}\right)\pmod{q}$
$=\sum_{i=1}^{k}B^{\tau}\cdot c_{i}\sigma_{i}\pmod{q}$
$=\sum_{i=1}^{k}c_{i}\mathbf{t}_{i}=\mathbf{t}.$
To sum up, the scheme satisfies the correctness with overwhelming probability. ∎
4 Security and privacy
In the field of provable security, the security of signature algorithms is usually related to hard computational problems. Typically, such proofs describe a reduction from a certain (assumed) hard computational problem to breaking the security properties defined by the signature system. Suppose there exists an adversary $\mathcal{A}$ who can break a signature scheme with a non-negligible advantage $\epsilon$ within polynomial time $t$. We construct a simulator $\mathcal{C}$ to solve a difficult computational problem with an advantage of $\epsilon^{\prime}$ within time $t^{\prime}$, where
$t^{\prime}=t+T,\quad\epsilon^{\prime}=\frac{\epsilon}{L}.$
If $(t^{\prime},\epsilon^{\prime})\approx(t,\epsilon)$, or more specifically, if $T=o(t)$ and $L$ is a constant or relatively small, then we say that this reduction process is a tight reduction, and we say that the signature scheme is tightly secure under the assumption that the computational problem is hard. This requires that the tight security of the signature algorithm be independent of the attacker’s parameters, such as the attacker’s number of queries or its security parameter.
In essence, a tight reduction means that a successful attack algorithm can be transformed into an efficient algorithm for solving difficult computational problems without significantly increasing the running time and or significantly reducing the success probability. Therefore, the tightness of the reduction determines the strength of the security guarantee provided by the security proof: a non-tight reduction provides a weaker security guarantee than a tight reduction. Therefore, it is highly desirable for a signature system to have a tight security reduction.
We will now prove that our linearly homomorphic signature scheme is tightly secure under the EUF-CMA security model in the standard model.
Theorem 10.
Suppose $q=\mathrm{poly}(n)$, $k=\mathrm{poly}(n)$, $q\geq(nk)^{2}$, $h=\lfloor\frac{n}{6\log q}\rfloor$, $V=\sqrt{2nk\log q}\cdot\log n$, $\beta=2k\cdot V\cdot\sqrt{2kn}$. If the $\textbf{SIS}_{q,\beta}$ problem is hard, then the above linearly homomorphic signature scheme is tightly secure under the EUF-CMA security model.
More specifically, let $\mathcal{A}$ be a polynomial-time adversary that makes at most $q_{s}$ signature queries to the simulator $\mathcal{C}$. If the adversary $\mathcal{A}$ breaks the above signature scheme with advantage $\epsilon$ within time $t$, then the simulator can construct a polynomial-time algorithm $\mathcal{B}$ that within $t+\mathcal{O}(q_{s}T_{\textbf{Sign}}+kT_{\textbf{SampleDom}})$ time outputs a solution to the $\textbf{SIS}_{q,\beta}$ problem with nearly the same advantage $\epsilon-\mathrm{negl}(n)$, where $T_{\textbf{Sign}}$ represents the time spent in generating a signature and $T_{\textbf{SampleDom}}$ represents the time for one execution of the sampling algorithm SampleDom.
Proof.
Assume there exists a polynomial-time adversary $\mathcal{A}$ that breaks the aforementioned signature scheme with advantage $\epsilon$ within time $t$. We will now prove that a simulator $\mathcal{C}$ can be constructed, which utilizes the adversary $\mathcal{A}$’s ability to forge signatures to construct an algorithm $\mathcal{B}$ that can output a solution to the $\textbf{SIS}_{q,\beta}$ problem.
Uniformly and randomly select a matrix $A\in Z_{q}^{h\times n}$ as the problem instance of the SIS problem. It is noted that, with overwhelming probability, matrix $A$ is a row-full-rank matrix. Now the simulator $\mathcal{C}$ generates the public key in the following way:
  1. (1)
    The simulator $\mathcal{C}$ runs the algorithm $\textbf{Setup}(1^{n},k)$ to obtain the public parameter $pp=(q,h,V)$.
  2. (2)
    Sample $k$ vectors using the algorithm $\textbf{SampleDom}(1^{n},s)$, where $s=\frac{V}{\sqrt{k/2}}\geq\omega(\sqrt{\log n})$:
    $\mathbf{x}_{i}\stackrel{{\scriptstyle\$}}{{\leftarrow}}\textbf{SampleDom}(1^{n},s),\quad\text{for}\quad i=1,...,k,$
    and then compute $\alpha_{1}=A\cdot\mathbf{x}_{1}(\mathrm{mod}q),...,\alpha_{k}=A\cdot\mathbf{x}_{k}(\mathrm{mod}q)$.
The simulator $\mathcal{C}$ outputs the public key $pk=(A,\alpha_{1},...,\alpha_{k})$ to adversary $\mathcal{A}$.
After receiving the public key $pk=(A,\alpha_{1},...,\alpha_{k})$, the adversary $\mathcal{A}$ selects a sequence of data sets $\mathbf{m}_{1},...,\mathbf{m}_{q_{s}}$ as signature queries, where $\overrightarrow{\mathbf{m}}_{i}=(\mathbf{m}_{i1},...,\mathbf{m}_{ik})\in(\mathcal{M})^{k}$. For each $i$ ($i=1,...,q_{s}$), the simulator $\mathcal{C}$ choose $\tau_{i}$ uniformly from $\{0,1\}^{n}$ and gives to $\mathcal{A}$ the tag $\tau_{i}$. The steps for the simulator to return the signature queries to the adversary $\mathcal{A}$ are as follows:
  1. (1)
    Take $\tau_{i}\in\{\tau_{1},...,\tau_{q_{s}}\}$, $\mathbf{m}_{ij}\in\{\mathbf{m}_{i1},...,\mathbf{m}_{ik}\}$, where $1\leq i\leq q_{s}$ and $1\leq j\leq k$.
  2. (2)
    Compute $(\mathbf{u}_{ij},\mathbf{v}_{ij}){\leftarrow}\textbf{Decompose}(\mathbf{m}_{ij})$, where $\mathbf{m_{ij}}=(m_{ij}^{(1)},...,m_{ij}^{(k)})=\mathbf{u}_{ij}+\mathbf{v}_{ij}$, $\mathbf{u}_{ij}=(u_{ij}^{(1)},...,u_{ij}^{(k)})\in\mathcal{X}$, $\mathbf{v}_{ij}=(v_{ij}^{(1)},...,v_{ij}^{(k)})\in\mathcal{Y}$ or $\mathcal{Z}$.
  3. (3)
    Compute $H_{\tau_{i}}=\mathrm{diag}\{2\tau_{i1}-1,2\tau_{i2}-1,...,2\tau_{in}-1\}$
  4. (4)
    Compute $\sigma(\mathbf{u}_{ij})=H_{\tau_{i}}\cdot\sum_{l=1}^{k}u_{ij}^{(l)}\mathbf{x}_{l}$, and $\sigma(\mathbf{v}_{ij})=H_{\tau_{i}}\cdot\sum_{l=1}^{k}v_{ij}^{(l)}\mathbf{x}_{l}$. According to Lemma 9, Lemma 10 and Corollary 7, $\sigma(\mathbf{u}_{ij})$ and $\sigma(\mathbf{v}_{ij})$ follow distribution $\mathcal{D}_{\mathbb{Z}^{n},V}$ within a negligible statistical distance.
  5. (5)
    Compute $\sigma_{ij}=\sigma(\mathbf{u}_{ij})+\sigma(\mathbf{v}_{ij}).$
The simulator $\mathcal{C}$ output the signature $\sigma_{ij}$.
Next, we prove that the output of the simulator is the same as the distribution in the real signature scheme (within a negligible statistical distance).
In the real scheme, the public-key matrix $A$ is generated by TrapGen and its distribution is statistically close to the uniform distribution over $\mathbb{Z}_{q}^{n\times m}$. In the simulation, $A$ is sampled uniformly at random from $\mathbb{Z}_{q}^{n\times m}$. Thus, the two distributions are statistically indistinguishable.
Secondly, $\alpha_{1},\alpha_{2},...,\alpha_{k}$ is also indistinguishable from the real scheme. This is because, according to Lemma 4, $\alpha_{j}=A\cdot\mathbf{x}_{j}(\mathrm{\mod}q)$ $(j=1,...,k)$ is statistically close to being uniformly distributed.
Finally, we demonstrate that the simulated signature is also statistically indistinguishable from the real signature. Since $\mathbf{x}_{j}\stackrel{{\scriptstyle\$}}{{\leftarrow}}\textbf{SampleDom}(1^{n},s)$ and $s=\frac{V}{\sqrt{k/2}}\geq\omega(\sqrt{\log n})$. Thus, by Lemma 5, $\parallel\mathbf{x}_{j}\parallel\leq\frac{V}{\sqrt{k/2}}\sqrt{n}\leq V\sqrt{n}$ holds with overwhelming probability. Therefore,
$\parallel\sigma(\mathbf{u}_{ij})\parallel=\parallel H_{\tau_{i}}\cdot\sum_{l=1}^{k}\mathbf{m}_{ij}^{(l)}\mathbf{x}_{l}\parallel=\parallel\sum_{l=1}^{k}\mathbf{m}_{ij}^{(l)}\mathbf{x}_{l}\parallel\leq kV\cdot\sqrt{n}.$
Similarly, $\parallel\sigma(\mathbf{v}_{ij})\parallel\leq kV\cdot\sqrt{n}$. Thus $\parallel\sigma_{ij}\parallel=\parallel\sigma(\mathbf{u}_{ij})+\sigma(\mathbf{v}_{ij})\parallel\leq 2kV\cdot\sqrt{n}\leq kV\cdot\sqrt{2kn}$.
In addition:
$AH_{\tau_{i}}^{\top}\cdot\sigma(\mathbf{u}_{ij})$ $=A\cdot H_{\tau_{i}}^{\top}H_{\tau_{i}}\cdot\sum_{l=1}^{k}u_{ij}^{(l)}\mathbf{x}_{l}=A\cdot\sum_{l=1}^{k}u_{ij}^{(l)}\mathbf{x}_{l}$ (2)
$=\sum_{l=1}^{k}u_{ij}^{(l)}A\cdot\mathbf{x}_{l}=\sum_{l=1}^{k}u_{ij}^{(l)}\alpha_{l}=\mathbf{t}(\mathbf{u}_{ij})(\mathrm{mod}q)$
Since $\sigma(\mathbf{u}_{ij})$ statistically follows the distribution $\mathcal{D}_{\mathbb{Z}^{n},V}$, but also satisfies the condition $\sigma(\mathbf{u}_{ij})\in\Lambda_{q}^{\mathbf{t}(\mathbf{u}_{ij})}(B^{\tau_{i}})$, where $B^{\tau_{i}}=AH_{\tau_{i}}^{\top}$. Therefore, according to Lemma 10, $\sigma(\mathbf{u}_{ij})$ statistically follows the distribution $\mathcal{D}_{\Lambda_{q}^{\mathbf{t}(\mathbf{u}_{ij})}(B^{\tau_{i}}),V}$. Similarly, $\sigma(\mathbf{v}_{ij})$ statistically follows the distribution $\mathcal{D}_{\Lambda_{q}^{\mathbf{t}(\mathbf{v}_{ij})}(B^{\tau_{i}}),V}$.
The adversary can also verify as follows:
$B^{\tau_{i}}\cdot\sigma_{ij}$ $=B^{\tau_{i}}\cdot(\sigma(\mathbf{u}_{ij})+\sigma(\mathbf{v}_{ij}))=B^{\tau_{i}}\cdot\sigma(\mathbf{u}_{ij})+B^{\tau_{i}}\cdot\sigma(\mathbf{v}_{ij})$ (3)
$=AH_{\tau_{i}}^{\top}\cdot(H_{\tau_{i}}\cdot\sum_{l=1}^{k}u_{ij}^{(l)}\mathbf{x}_{l})+AH_{\tau_{i}}^{\top}\cdot(H_{\tau_{i}}\cdot\sum_{l=1}^{k}v_{ij}^{(l)}\mathbf{x}_{l})$
$=A\cdot\sum_{l=1}^{k}u_{ij}^{(l)}\mathbf{x}_{l}+A\cdot\sum_{l=1}^{k}v_{ij}^{(l)}\mathbf{x}_{l}$
$=\sum_{l=1}^{k}u_{ij}^{(l)}A\cdot\mathbf{x}_{l}+\sum_{l=1}^{k}V_{ij}^{(l)}A\cdot\mathbf{x}_{l}$
$=\sum_{l=1}^{k}u_{ij}^{(l)}\alpha_{l}+\sum_{l=1}^{k}v_{ij}^{(l)}\alpha_{l}$
$=\sum_{l=1}^{k}(u_{ij}^{(l)}+v_{ij}^{(l)})\alpha_{l}=\mathbf{t}_{ij}.$
Hence, the simulated signature and the real signature are also statistically indistinguishable.
After the adversary $\mathcal{A}$ receives all the queried signatures, it will generate a valid forgery $(\tau^{*},\mathbf{m}^{*},\sigma^{*},f)$, where $\mathbf{m}^{*}=(m_{1}^{*},...,m_{k}^{*})$, $\langle f\rangle=(c_{1},...,c_{k})$ , such that
$\textbf{Verify}(pk,\tau^{*},\mathbf{m}^{*},\sigma^{*},f)=1.$
Whether it is a I-type forgery or II-type forgery, the following equation holds:
$B^{\tau^{*}}\cdot\sigma^{*}(\mathrm{mod}q)=AH_{\tau^{*}}^{\top}\cdot\sigma^{*}=\sum_{i=1}^{k}m_{i}^{\ast}\alpha_{i}=\sum_{i=1}^{k}m_{i}^{\ast}A\cdot\mathbf{x}_{i}=A\cdot\mathbf{x}^{\ast}(\mathrm{mod}q),$
where $\mathbf{x}^{\ast}=\sum_{i}^{k}m_{i}^{\ast}\mathbf{x}_{i}$. By Lemma 5, $\parallel\mathbf{x}_{i}\parallel\leq s\cdot\sqrt{n}=\frac{V}{\sqrt{k/2}}\cdot\sqrt{n}$ holds with overwhelming probability, so $\parallel\mathbf{x}^{\ast}\parallel\leq V\cdot\sqrt{2kn}\leq k\cdot V\cdot\sqrt{2kn}$. By verification condition (1), $\parallel H_{\tau^{*}}^{\top}\cdot\sigma^{*}\parallel=\parallel\sigma^{*}\parallel\leq k\cdot V\cdot\sqrt{2kn}$, thus $\parallel H_{\tau^{*}}^{\top}\cdot\sigma^{*}-\mathbf{x}^{\ast}\parallel\leq 2k\cdot V\cdot\sqrt{2kn}$.
Let the process of outputting $H_{\tau^{*}}^{\top}\cdot\sigma^{*}-\mathbf{x}^{\ast}$ be denoted as algorithm $\mathcal{B}$. If $H_{\tau^{*}}^{\top}\cdot\sigma^{*}-\mathbf{x}^{\ast}\neq 0$, then the simulator outputs $H_{\tau^{*}}^{\top}\cdot\sigma^{*}-\mathbf{x}^{\ast}$ as a solution to the $\textbf{SIS}_{q,\beta}$, where $\beta=2k\cdot V\cdot\sqrt{2kn}$.
Finally, we discuss the probability of $H_{\tau^{*}}^{\top}\cdot\sigma^{*}=\mathbf{x}^{\ast}$. According to the properties of the sampling algorithm, that is, the preimages have a conditional min-entropy of $\omega(\log n)$. Therefore
$\mathrm{Pr}[H_{\tau^{*}}^{\top}\cdot\sigma^{*}=\mathbf{x}^{\ast}]\leq 2^{-\omega(\log n)}\leq\mathrm{negl}(n).$
Consequently, Algorithm B will output a solution to the $\textbf{SIS}_{q,\beta}$ problem with an advantage of $\epsilon-\mathrm{negl}(n).$
Worst-case to average-case reduction. Based on Theorem 1 from the reference [32], when $q\geq\beta\cdot\omega(\sqrt{n\log n})$ holds, the computational hardness of the $\textbf{SIS}_{q,\beta}$ problem is equivalent to that of approximating the SIVP problem in the worst-case setting with an approximation factor of $\beta\cdot\tilde{O}(\sqrt{n})$. In the Setup algorithm, the requirement $q\geq(nk)^{2}$ serves to ensure that $q$ has a large enough value for Theorem 1 to be applicable.
From the proof of Theorem 10, it can be observed that neither a I-type forgery nor a II-type forgery requires the adversary to specify a target tag $\tau_{g}$ in advance. Therefore, our signature scheme is also tightly secure under the U-ST-SCMA security model.
Theorem 11.
Suppose $q=\mathrm{poly}(n)$, $k=\mathrm{poly}(n)$, $q\geq(nk)^{2}$, $h=\lfloor\frac{n}{6\log q}\rfloor$, $V=\sqrt{2nk\log q}\cdot\log n$, $\beta=2k\cdot V\cdot\sqrt{2kn}$. If the $\textbf{SIS}_{q,\beta}$ problem is hard, then the above linearly homomorphic signature scheme is tightly secure under the U-ST-SCMA security model.
Finally, we prove that our linearly homomorphic signature scheme is weakly context hiding. We have the following theorem.
Theorem 12.
Assume that $V>k^{s/2}\cdot\omega(\sqrt{1+\log n})$, then the above linearly homomorphic signature scheme is weakly context hiding.
Proof.
The challenger runs KeyGen, generating a public-private key pair $(pk,sk)$, and sends it to the adversary. Upon receiving the public-private key pair, assume the adversary outputs $(\mathbf{m}_{0},\mathbf{m}_{1},f_{1},...,f_{s})$ to the challenger, where $\mathbf{m}_{0}=(\mathbf{m}_{01},...,\mathbf{m}_{0k}),\mathbf{m}_{1}=(\mathbf{m}_{11},...,\mathbf{m}_{1k})$ and satisfies $f_{i}(\mathbf{m}_{0})=f_{i}(\mathbf{m}_{1})$, $i=1,2,...,s.$
For convenience, we treat $\mathbf{m}_{b}$ as a matrix over $\mathbb{F}_{2}^{k\times k}$, where the $j$-th row corresponds to the $j$-th message in the data set $\mathbf{m}_{b}$. We denote $<f_{i}>=(c_{i1},...,c_{ik})$ and construct the $s\times k$ matrix $G$, where the $i$-th row is $<f_{i}>$.Therefore, we have
$G\mathbf{m}_{0}=G\mathbf{m}_{1}.$
In response, the challenger randomly selects a tag $\tau\in\{0,1\}^{n}$ and a random bit $b\in\{0,1\}$. We define $\sigma_{j}^{(b)}$ as the signature of the $j$-th message in the data set $\mathbf{m}_{b}$. We then construct a $k\times n$ matrix $\Sigma^{(b)}$, where the $j$-th row is $(\sigma_{j}^{(b)})^{\top}$. For each $i=1,2,...,s$, let $d_{i}^{(b)}=\textbf{Evaluate}(pk,\tau,f_{i},(\sigma_{1}^{(b)},...,\sigma_{k}^{(b)}))$. Then, we construct an $s\times n$ matrix $D^{(b)}$, where the $j$-th row is $(d_{j}^{(b)})^{\top}$. According to the definition of Evaluate, we have
$D^{(b)}=G\Sigma^{(b)}.$
For $b=0,1$, $\sigma_{i}^{(b)}$ follows the distribution $\mathcal{D}_{\Lambda^{\tau}+\mathbf{t}_{i}^{(b)},\sqrt{2}V}$, where $\mathbf{t}_{i}^{(b)}=\sum_{j=1}^{k}m_{bi}^{(j)}\alpha_{j}$, and $\mathbf{m}_{bi}=(m_{bi}^{(1)},...,m_{bi}^{(k)})$. By the definition of the signing algorithm, the signatures are independent of each other. Therefore, when $V$ is sufficiently large $(>\mu(G)\eta_{\varepsilon}(\Lambda^{\tau})$, by Theorem 5, the distributions of $D^{(0)}$ and $D^{(1)}$ are statistically close. Therefore, any polynomial-time adversary cannot distinguish between $D^{(0)}$ and $D^{(1)}$ with non-negligible probability and further guess the value of $b$. Now we only need to prove that
$V>k^{s/2}\cdot\omega(\sqrt{1+\log n})\geq\mu(G)\eta_{\varepsilon}(\Lambda^{\tau}).$
First, since the elements of the matrix $G$ are from the set $\{0,1\}$, then by the Hadamard inequality:
$\mu(G)=\sqrt{det(GG^{\top})}\leq\prod_{i=1}^{s}\parallel\mathbf{g}_{i}\parallel\leq k^{s/2}<V,$
where $\mathbf{g}_{i}$ represents the $i$-th row of the matrix $G$.
Secondly, According to Lemma 6, there exists a negligible $\epsilon$ such that
$\eta_{\epsilon}(\Lambda^{\tau})\leq\omega(\sqrt{1+\log n})<V$
holds with overwhelming probability. Therefore,
$V>k^{s/2}\cdot\omega(\sqrt{1+\log n})\geq\mu(G)\eta_{\varepsilon}(\Lambda^{\tau}).$
5 Conclusion
Constructing tightly secure short homomorphic signature schemes based on the hardness of the Short Integer Solution (SIS) problem in the standard model remains a longstanding open problem. This paper presents the first linearly homomorphic signature scheme with tight security based on the SIS assumption in the standard model. To achieve this, we introduce two key algorithms—NewBasis and Decompose—for generating shorter signatures and enabling tight security reductions. We conjecture that these algorithms may have potential applications in other cryptographic constructions.
Several unresolved challenges persist in current research: As noted in [21], designing more efficient signatures and encryption schemes (such as hierarchical encryption, attribute-based encryption, hierarchical signatures, homomorphic signatures supporting polynomial functions, and hierarchical fully homomorphic signatures) with tight or nearly tight security in the standard model remains an open issue. Furthermore, another promising direction involves constructing secure homomorphic signatures based on semigroups—a problem posed in [8] that remains unexplored.